A Complete Bibliography of Publications in Transactions of the American Fisheries

Society: 1930–1939

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

19 September 2022 Version 1.01

Title word cross-reference

1930 [Ano30a]. 1931 [Ano31a]. 1932 [Ano32g]. 1933 [Ano33g]. 1934 [Wes34]. 1935 [Wic35]. 1936 [Bel37b]. 1937 [Ano38h].

64th [Wes34]. 65th [Wic35]. 66th [Bel37b]. 67th [Ano38h].

Ability [WM32]. Abnormally [WM32]. Abstracts [Ano32a]. Abundance [Hil37]. Account [Smi38]. Accurate [Pre34b]. Acriflavine [Atk32a]. Activities [Pow31, Sc.32]. Adaptability [Moo32]. Additional [Wic38]. Adirondack [Fai34]. Adjustable [Nee35]. Affect [Joh34, Mot38a]. Affecting [Dym30]. after [Bel34a]. Agriculture [Age34a]. agua [Cur34]. agua-bonita [Cur34]. aid [Wer34]. Aim [Cle32]. Air [Pre35, Wol37a]. Alewife [Ode34]. Algae [M’G34]. along [VW38b]. Alosa [VW38b]. amended [Ano30c, Ano31c, Ano32c, Ano33c, Ano34b, Ano35b, Ano37b, Ano38b]. American [Ano30a, Ano31a, Ano32g, Ano33g, Ano38h, Bel37b, Wes34, Wic35, Ano35b, Ano37b, Ano38b, O’M32]. Americanized [Mar35]. Anabiosis [Bor32]. Analyses [Age34d]. Analysis [Esc35]. Anchor [Tid34]. Angler [Bel30, Cro37, Cla34]. Animal [Nur30]. Annual [Ano30a, Ano31a, Ano38h, Bel34b, Bel37b, Han37, Wes34, Wic35]. anomalum [Smi35c]. Aplites [Coo37]. Apparatus [Pre34b]. Aquatic [Cul30, Ols32, RW34, Spi34, SM31]. Aquicultural [Lan35a]. Aquiculture [Age34a, Hun34b]. Areas [Rod30]. arenaria [Mar30]. Arkansas [Ano31a, Bro31]. Arsenic [SM31, Wie30, WGS31]. Arsenite [Sur31a]. Artemia [Sea33]. Artificial [Age33, But37, Foe31, Foe32, Hil37, Smi37a, Wer34]. Artificially [Age34d]. Aspects [Mee34]. Association [Bel30]. Atlantic [BCF37, BH32, Bel34a, Bel34c, BK34, Bel35a, Bel37a, VW38b, WLA32, Whi34a]. Atmosphere [Wie31]. Attack [Hun30]. August [Ano30a, Ano38h].

Bacteria [Mar30]. Bacterial [Atk32a, Fis35, GS34]. Balanced [Age34b]. Baltimore [Ano32g]. Barren [Atk32b]. Bass [Ban33, Coo37, Dav30, Den31, Den34, Haw33, Hay32a, Hay33, Hog33, Hog34, How31, How32, Hub31, Jam30, Kin34, KR35, Kin37, Lan34a, Lan34b, Lan37a, Mar32, Mil32, Pow31, Raw38, Reg34, Riv37, Shi31, Sny32, Tes30, TV37b, Vio31b, VW38a, WGS31, WF33]. Bass-Brooding [Reg34]. Basses [Bro31]. Bay [Sto38, Ear32, Sch37a, TV37b, VW38a]. Bearing [RW34]. Beaver [Bra35]. Beaver-Trout [Bra35]. Behavior [Lan34a]. Benefits [Bel38b]. Better [Vio35]. Big [Nur30]. bilinearis [BCF37]. Biological [Roa33, Sch37a, Sch37b]. Biology [Fie32, Sur30b]. Black [Ban33, Den31, Den34, Haw33, Hog33, Hog34, How31, How32, Hub31, Mil32, Raw38, Shi31, Sny32, Tes30, Vio31b, WGS31, WF33, WE32]. Black-Headed [WE32]. Blackspotted [Lor30]. Bladder [Wol37a]. Blood [Joh34, Wol37b]. Blue [WF35]. Body [MTC^{+}31]. bonita [Cur34]. Book [Bos37]. Bottle [Dwy32]. Bottom [Nee34, Sur30b, Sur37]. Bower [Bow37]. Branch [She37]. Brazil [vIW35]. Breeding [Cat30, HE30, McN37, Shi34, Smi35c, WE32]. Brine [Sea33]. British [Mot32, Pri32, Cle30, Dym30, Rus37, Tes33]. Brome [Phi34, Spi34]. Brood [Lor30]. Brooding [Reg34]. Brook [Det30, Din34, Gre32, HE30, Haz32, HJ38, Lor33, Nee30, She37, SA32, TM33, Whi30, Whi34b, Lor35]. Brown [Gre32, Mar35, SA32]. Buckeye [EB35]. Bullfrog [Vio31a]. Bullheads [Atk31]. Burbot [Bra30]. Button [Van37a]. By-Laws [Ano30c, Ano31c, Ano32c, Ano33c, Ano35b, Ano37b, Ano38b, Ano34b]. By-Products [Don35].

C.C.C. [HTE33]. Calculating [Mot37]. Caledonia [Smi33]. California [CC33, Cro37, Nee38, Puj38, Sco32, Taf34, Wal35]. Callinectes [WF35]. Cambarus [Lan35b, Lan37b]. Campaign [Loc32]. Campostoma [Smi35c].

Can [Joh34]. Canada [Wes34, Cle32]. Canadian [BPR32, Bor34]. Cannibalism [Coo37]. carassii [Tid34]. Carbon [Sur35a]. Carp [Bur31, Tid34]. Cat [JSR33, Mob31]. Catch [Cla34, Cro37, Esc35]. Catches [CC33]. catesbiana [Vio31a]. Catostomus [Hun35]. caurinus [Sch35]. Cause [Age33, Bel34a]. Census [Esc37, Moo38, Smi37c]. Central [CT30, Kos37]. Certain [Elk37, Moo32, Ode32, Smi38, SM31, Van37a]. Certificate [Ano30b, Ano31b, Ano32b, Ano33a, Ano34a, Ano35a, Ano37a, Ano38a]. Changes [Mar32, Wie31]. Channel [JSR33, Mob31]. Characteristics [Fai34, Mar35]. Chautauqua [Moo38]. Cheap [Nee35, Tay35]. Chemical [Dom35, Fai34, Fis33, GS34, MTC^{+}31, Roa33]. Chequamegon [Elk37]. Chesapeake [Ear32, TV37b, VW38a]. Chub [Sch35]. Circular [CT30, SA32, Sur33a]. Cisco [Van30]. Ciscoes [Sto38]. City [Ano38h]. Clam [Mar30]. Clean [Loc32, Loc34]. Clots [Wol37b]. clupeaformis [HD34]. Coal [Pau38]. Coarse [Cat34, Cle34, Wie34]. Coast [Bel37a, Duf32, Pri32, VW38b]. Coastal [Hac34, Tes33]. Coasts [WLA32]. Collected [Sco32]. Collecting [CC33]. Columbia [Cle30, Dym30, Mot32, Pri32, Tes33]. Columbus [Ano33g]. Combating [Ken32]. Comments [Sco32]. Commerce [Fie37]. Commercial [Gow32a, HD33, Hog33, Smi37a, Van32, Van34]. commersonii [Hun35]. Committees [Ano33b, Ano34g, Ano34f, Ano35g, Ano35f, Ano37g, Ano37f, Ano38g, Ano38f]. Common [WLA32]. Comparative [Atk32b]. Comparison [CF34, Foe31, HE37]. Composition [BCF37, MTC^{+}31]. Computing [Lan34b]. Concentration [Wie31]. Concentrations [SM31, Wie30, WM32]. concerning [Bor34, Van35, Whi34a]. Conditions [Baj35, HT35, Moo32]. Conduct [Gow32a]. Conservation [Bel34d, Cle30, Her32, Osb33, Pri32, RA34, Wic31]. Constitution [Ano30c, Ano31c, Ano32c, Ano33c, Ano34b, Ano35b, Ano37b, Ano38b]. Consumption [Baj35, Mar32, WF33]. Container [Loc30]. Content [Smi34, WGS31]. Control [Bur31, Hun34b, Wol35]. Controlling [LE31, Pen30, Sur31a]. Convenient [Loc30]. Conveyor [Bow37]. Cooperation [Bel35b, Bel38b, Ear37]. Cooperative [Mar33b, Mil37]. Copepod [Sav35a]. Copper [Cat34, Smi35b]. Coregonus [HD34]. cornutis [Nur31]. Cost [Deu34]. Costiasis [Sav35b]. Cottidae [Kos37]. Cottonseed [Deu35]. Cottonwood [Cur34]. Count [GS34]. Counting [Pre34b]. County [Ban37a, Mar33b, TL34]. Course [Cle31]. Crab [WF35]. Crappie [Han37]. Crayfish [Lan35b, Lan37b]. Creel [Esc37, Moo38, Smi37c]. Cristivomer [VD38]. Critical [Dym30]. Criticism [Pre34a]. Crude [Wie35]. Crustacea [Ewe33]. Cuban [Riv37]. Cultural [Cle31, Hay32b, Sc.32, Sur31b]. Culture [Ald35, Bos37, But30, Cle32, Cle34, Coo32, Cul32, Dav30, Hay32a, Hub33, Hun38, Lan32, Lan33, M’G32, Mac30b, MW34, McL30, Mob31, Mot35, Pre34a, Vio31a, Vio35]. Cutthroat [HM33, Pau38]. Cyclochaetosis [Ric38].

Daily [Baj35]. Dam [Foe30]. Dams [Har32, Ric34]. Daphnia [ES34, Smi37b]. Data [Tra32]. Date [Han37, Ano30c, Ano31c, Ano32c, Ano33c, Ano34b]. Dauphin [Cun35b]. Decrease [McK30]. Deductions [Van35]. Deep [Wes32]. Defined [Wer34]. Density [She37]. Depletion [Cul30, HJ38]. Depredations [LS37]. Depressions [Wie34]. Depth [Ode32]. Destruction [Atk31, Bea37]. Determine [Cob33]. Determining [Bel34b, CF34, GS34, Wic33a]. Development [Sur35a]. Developments [Loc32]. Devices [Coo32]. Diet [Age34b, Dav32]. Diets [Deu35]. Dikerogammarus [Smi33]. Dioxide [Sur35a]. Disappearance [Van30]. Discocotyle [LE31]. Disease [Age33, Atk32a, BM35, Fis34, Fis35, Wol37a]. Disinfection [Fis33]. Dispersal [Mee35, Van37b]. Dissolved [Smi34, WM32, Wie34]. Distribution [Ode32, Pen30, Phi34]. District [Mot32]. Diversions [Bel31]. Division [Wic31]. do [Age34c, TH35]. Does [Mot38a]. doing [Ear30]. Dominance [Dea33]. Dominating [Spi34]. Dried [Deu35]. Drought [Jam34, LeC30]. Dry [Dav32, Smi37c]. Ducks [Bea37]. during [Coo37].

Early [BPR32]. Eastern [HD34, Haz32, Whi30]. Ecological [Mee34, Ode34]. Economic [Hay32a, RA34, Sch37a]. Edward [Ric37, Whi30]. Effect [Din34, Dom35, Jam34, LeC30, Mar30, Pau38, Sav37, Wie35]. Effects [Fie30, M’G32, Mar33a, Ric34, Sur35a, Tra33]. Efficient [Lan31]. Effluent [RA34]. Egg [Deu35, Foe34, Gre32]. Eggs [Atk31, Atk32a, Emb34, Sur35a, Wic33a]. Election [Ano30d, Ano31d, Ano32d, Ano33d, Ano34c, Ano35c, Ano37c, Ano38c]. Electricity [Bur31]. Eleven [Dom35]. Elimination [Cat34]. Employed [Sur30b]. Enclosures [Raw38]. Encountered [Kin37]. End [Boe38, Ewe33]. Enterprise [Mil37]. Entire [MTC^{+}31]. Equipment [Cob32, Reg34]. Eradicating [Smi35b]. Erie [Boe38, Dea33, Ewe33, Van30, Van32, WT33]. Errors [Mot37]. Esox [McN37, McN37]. Estimate [Baj35]. European [Mar35]. Evermann [Nee38]. Evidence [Tar37]. Exceptionally [Haz35]. Excess [Wie34]. Exemplified [Rad33]. Experience [McL30]. Experiences [Hub33]. Experimental [DF38, Esc38, Lor32, Pre35, Smi32, Tar37]. Experiments [Bor32, Bro31, Cro30, Det30, Hog33, How32, Sur35b, Taf34, Van32]. Exploded [Van35]. Exploring [Wes38]. Exposure [Wie30, Wie31, WM32]. Extent [Mar32]. Eyed [Foe34]. Eyed-Egg [Foe34].

Fact [Din35]. Factor [Dym30, Haz33, Hew37, M’G34]. Factors [Hun34a, Pen30, Pow38, Tar38]. Factory [Fie37]. Facts [Whi34a]. Facultative [Sur30b]. Failures [Sur31b]. Fall [BK34]. Fall-Run [BK34]. Farming [Pry32, Rad33, Rad38]. Fasciatus [Sur35b]. Fate [Mar34]. Fauna [Sur30b, Sur37]. Fecundity [Sto38]. Federal [Bel35b, Den31]. Feeding [BH32, Det30, DF38, FS35, Lor33, Ric30, Sur35b, Tay35, Wal35]. Fertility [Wic33a]. Fertilization [Smi32]. Fertilized [Smi34, Smi37b]. Fertilizers [Hog33, Mee33, Mee34, Wie34]. Fertilizing [Mee35]. Fields [Wes38, WBF34]. Fife [Esc37]. Fight [Mil32]. Fingerling [SA32, WF33]. Fingerlings [CT30, Det30, Hog33, Pre34b, WGS31, Wol37a]. First [Ano31a, Kin34]. Fish [Ait37, Ald35, Baj32a, Baj32b, Baj35, Ban37a, Ban37b, Bea37, Bel32, Ber38, Boe38, Bos37, Bow37, But30, Cat34, Cla33, Cle31, Cle32, Cle34, CF34, Coo32, Coo35, Cul32, Deu35, Dom35, Ear37, Elk37, Esc35, FS35, Har32, Hay32a, Hay32b, Her32, Hub33, HT35, Hub37, Hun38, Jam34, Joh34, Lan33, Lan35b, Loc30, Lor34, M’G32, Mac30b, Mar33b, Mar34, McK30, McL30, Mob31, Mot35, Mot37, NB33, Nee35, Nes33, Ode32, Osb33, Par34, Pre34a, RA34, RW34, Sc.32, Sch35, Smi35a, Smi35b, Smi37a, Smi38, Sur31b, Sur31a, SO37, Tar38, Tra33, Vio35, VW38a, War31, Wat32, Wic32, Wic33b, Wic38, Wie30, Wie31, WM32, Wie34, Wie35, Wol35, Haz35]. Fisheries [Ano30a, Ano31a, Ano32g, Ano33g, Ano35b, Ano37b, Ano38b, Ano38h, Baj30, Bel37b, Cah37, Cul34, Dwy32, Ear32, Gor33, Gow32a, Hac34, Har30, HD33, Lan38, Mén38, O’M32, Puj38, Sch37b, Sco32, Tru32, Van32, Van34, Van35, Van37c, Wes32, Wes34, Wes38, Wic33c, Wic35, vIW35]. Fishery [Ano32a, Bel33, Bel35b, Bel38b, Ell37, Gow32b, Hub30, Hun30, Rus37, Wic31]. Fishes [Age33, Bor32, Bor34, Ewe33, EB35, Lan31, Mad37, Mar33a, Nee33, Nur30, Ols32, Pow38, Riv37, Ros31, Sea33, Sur30a]. Fishing [Baj30, Fra38, HE37, Mot38a, TV37a, Wat32]. Fishway [Web31]. Fishways [Bru30]. Fleas [NB33]. Floating [Reg34]. Florida [War31]. floridana [Mar32, WF33]. Foes [Kin37]. Follow [Moo31]. Follow-up [Moo31]. Fontinalis [Whi30, Whi34b, Haz32, Ric38]. Food [Baj35, Boe38, Coo37, Deu34, Don35, Dou35, Ewe33, EB35, Foe32, HM33, Kos37, Lor33, Mar32, McN37, Nee30, Nee33, NB33, Nur30, Sea33, Smi37b, Sur30a, Sur35b, Tar38, VD38, War31]. Foods [Age34d, Dav32, Dom35, FS35, Nee34]. Forage [Coo35, ES34, Hub33, Mar34, Sch35]. Foragers [Lor34]. Foreign [Ano32a]. Forest [Elk37]. Forests [Dav35]. Formation [Han37]. Forward [Hub37]. Four [Emb34]. Freezing [Bir32]. Fresh [Ell35, Mar33a, Ric32, Rus37, Tit30, Wie35]. Fresh-Water [Ric32, Tit30, Wie35]. Freshness [GS34]. Freshwater [Ros31]. Fry [Atk32a, CT30, Foe34, Pre34b, Rob37]. Full [Mot38a]. Fungisitosis [Age33]. Furnace [Lor35]. Further [HE30, Lan37b]. Furunculosis [Duf32, SK31]. Future [Cul34, Fie30, Gor33, Hun38].

Gambusia [War31]. Game [Baj32b, Esc35, Lor34, Mot35]. Game-Fish [Esc35]. Gammarus [Pen30, Sur35b]. Garfish [Bur31]. Gaspé [TL34]. Geer [Smi37b]. General [Tra33]. Gill [Fis35, LE31]. Gills [Wol37b]. Girard [Lan35b, Lan37b]. Golden [Cur34]. Goldfish [JSR33, Tid34]. Government [Ear37]. Grader [Nee35]. Grading [Bel32]. Grand [Bel37b]. Gravel [Rob37, Sav37]. Gravid [Van37a]. Graylings [Lor32]. Great [HD33, HP37, Smi37a, Van35, Van37b, Van37c, Wes32]. Green [Sch37a]. Grey [Atk31]. Grounds [BH32, Ken32]. Group [Esc38]. Growth [BPR32, Bel32, Bel34b, Bel35a, Coo37, Dav34a, Dea33, Din34, Gre33, HD34, Lin30, She37, Sto38]. Gulf [WLA32].

H [Hew37]. Habits [Bel34c, Coo37, Gre32, Lan35b, Lan37b, McN37, Ric30, Sch35, Smi35c, Sto38, Tes30, WE32]. Hackettstown [Hay32a]. Haddock [GS34]. Halibut [Bel33]. Hampshire [Smi37c]. Harvest [Lor35]. Hatcheries [BPR32, Ber38, Deu34, Lan31, Wal35]. Hatchery [Ald35, BM35, FS35, Kin34, Lor32, Lor34, M’G34, HE30, Hay32a, Hay32b, KR35]. Hatching [Rob37, Tay35]. Headed [WE32]. Health [Lin30]. her [Ear30]. Heredity [Dav31, Dav34a, Din34]. Herein [Wer34]. Herring [HT34, Tes33]. High [Bel34a, Har32, WM32]. Highlands [HD34]. Hill [Nee33]. History [Haz32, Ode34, WLA32]. Holbrooki [War31]. Hole [Bor32]. Hot [Ano31a]. Hubbs [Ban33, How31, Vio31b]. Huro [Mar32, WF33]. Huron [Hil37]. Hydrogenion [Wie31]. Hydrographic [Dwy32, Gow32a].

Ice [Ait37]. Ictaluras [Mob31]. Identification [Hew37]. II [Mee34]. Immature [Her32]. Importance [Gow32a, Gow32b, Par34, Sch37a, TV37a]. Important [Osb33]. Impounded [Ell37, Moo31, Moo32, Roa33, Wic32, WR37]. Improved [Bel34b, Cob32]. Improvement [Ait35, Dav34b, Gre35, HTE33, Tar31, Tar37]. Incorporation [Ano30b, Ano31b, Ano32b, Ano33a, Ano34a, Ano35a, Ano37a, Ano38a]. Increase [Ear37, Hil37]. Increasing [Cla33]. Incubation [Emb34]. Indiana [Shi31]. Industries [Bir32]. Industry [Baj30, Fra38, Rad33, Van37a]. Inexpensive [Age34b]. Infection [Lan37a, Ric38, Sav35a]. Infestations [Tid34]. Influence [Bir32, Dav31, Mar32]. Influencing [Hun34a, Tar38]. Ingestion [Mar30]. Inland [Vio35]. Institution [Bor32]. Instruction [Cle31]. Intensively [ES34]. Internal [Nes33]. International [Bel38b]. Interrelationships [Ols32]. Intestinal [Age33]. Introduced [Riv37]. Introduction [Bor34, Nee38]. Investigating [HT35]. Investigation [Bel33, HD33]. Investigations [Baj30, Bel35b, Coo35, Hea30, Ric32, SK31, TV37b, WT33, vIW35]. Involved [Pow38]. Iowa [Ait34, Ait35, Ait37]. Irondequoit [Sto38]. Irradiation [Cro30]. Island [Whi30]. Isles [Rus37].

Jersey [HE30, Hay32b]. Jesse [Smi35a]. Job [Gor33]. Jordan [Cur34]. Justified [Van35].

Kamloops [Mot32]. Kenora [McK30]. Kentucky [How31]. Kettle [Ait34]. Key [Vio35]. King [Don35]. Klamath [Dou35].

Laboratory [CF34, Har30]. Lacépède [WGS31]. Lake [Ait37, Esc35, Haz35, HD34, Joh34, Mil37, Moo38, Ode34, Raw32, Smi35b, Spi34, Sur33c, VD38, Wol37a, WT33, Boe38, Cun35a, Cun35b, Dea33, Esc37, Ewe33, EB35, Nur30, Ric37, Smi35a, Sny32, Van30, Van32, VD38]. Lakes [Atk32b, Baj32a, Dou35, Dym30, Elk37, Esc38, Fai34, HE37, Ode32, Ric34, Smi38, TL34, Cur34, Dom35, HD33, Hil37, HP37, Phi34, Smi37a, Van35, Van37b, Van37c, Wes32]. Lamprey [HP37]. Large [Mar32, WGS31]. Large-Mouth [Mar32, WGS31]. Largemouth [Ban33, Coo37, Hay33, Hog33, Hog34, Riv37, WF33]. Latreille [WF35]. Law [Den31, Wat32]. Lawrence [Fol32]. Laws [Ano30c, Ano31c, Ano32c, Ano33c, Ano35b, Ano37b, Ano38b, Ano34b]. Lawyer [VD38]. Learned [Din32]. Legislation [Den34, Van32]. Lernaea [Tid34]. Lethal [SM31]. Life [Haz32, Jam34, McK30, Ode34, Ols32, RW34, Tra33, WLA32, Wic32]. Limited [Cob31]. Limiting [Haz33]. Limnological [WT33]. lineatus [VW38a]. Ling [Bra30]. List [Ano30d, Ano31d, Ano32d, Ano33d, Ano34c, Ano35c, Ano37c, Ano38c]. Literature [Ano32a]. Live [Sea33]. Lobster [Cob32]. Logically [Van35]. Lonoke [Hog34]. Looking [Hub37]. Loss [Mot38b]. Lota [VD38]. Louisiana [Gow32a]. Low [Haz33]. Lower [Puj38]. lucius [McN37]. Luxilus [Nur31].

maculosa [VD38]. Made [Bor32, Cob33, Rod30]. Madison [Dom35]. Making [Cob31]. Man [Hac34]. Management [Ber38, Bra35, Dav35, Esc38, Gor33, Hub37, Lan31, Lan38, MGG^{+}34, Sur33a, Van37c, Wes38]. Manitoba [Baj32b]. Manitou [Spi34, Phi34]. Margaree [BC38]. Marginal [LH35]. Marine [CC33]. Marketing [Bra30]. Marking [Nes33]. Maryland [Ano32g, Ear30, Pow31, Tru32, TV37a]. Maskinonge [MW34]. Massachusetts [Ken32]. Master [Vio35]. Maturity [Dea33, Sur33c]. Mayflies [Nee33]. Meal [Deu35]. Meals [CF34]. Means [Bur31, Nes33]. Measure [Cla34]. Measurements [Bel34b]. Measures [Cul30]. Measuring [Cro37, Ell35]. Meat [Deu35]. Mechanical [Coo32]. Meeting [Ano30f, Ano30a, Ano31f, Ano31a, Ano32f, Ano33f, Ano34e, Ano35e, Ano37e, Ano38e, Ano38h, Bel37b, Wes34, Wic35]. Members [Ano30d, Ano31d, Ano32d, Ano33d, Ano34c, Ano35c, Ano37c, Ano38c]. Membership [Ano30d, Ano31d, Ano32d, Ano33d, Ano34c, Ano35c, Ano37c, Ano38c]. Menace [Ken32]. Merganser [LS37]. Merluccius [BCF37]. Mesh [Van32]. Metabolism [Cla33]. Method [Bel32, CC33, Har32, Nes33, Pre34b, RB34, Sur30b, Sur30a]. Methods [Bel34b, HD33, Lan34b, Pry32]. Mexico [Ano38h, Ber38, Fra38, Mén38, Nee38]. Michigan [Bel37b, Bra35, Coo32, Coo35, Coo37, Esc35, Esc38, Gre33, HE37, Hil37, Hub30, HTE33, LS37, Mil37, Sal33, She37, Tar31, Tar37, Van37a, VD38]. Micropterus [Ban33, How31, Vio31b, WGS31]. Migration [HJ38, Mot33, She37]. Migrations [Cob33]. Migratory [VW38a]. Mile [Sur33b, Sur37]. Milk [Deu35]. Minnesota [Sur31b, SO37]. Minnow [Smi35c, WE32]. Minnows [Moo32, Rad31]. Mississippi [Cul30, Wic30, Cul34]. Missouri [Ros31]. Mitchell [Haz32]. Mitchill [HD34, Hil37, RB34, Ric38, Van37b]. Moisie [Bel38a]. Monroe [Mar33b]. Montana [Lor32]. Montreal [Wes34]. Moon [Mot38a]. Moore [Ric35]. mordax [RB34, Van37b]. Mortalities [M’G34]. Mortality [Bel34a, Pow38]. Mosquito [War31]. Most [Cob31]. Mouth [Mar32, Vio31b, WGS31]. Mouthed [Jam30, Tes30]. Movement [Ait37]. Municipal [Ald35]. Mussels [Van37a]. Mya [Mar30]. Mylocheilus [Sch35].

N [Sny32]. namaycush [VD38, VD38]. National [Dav35, Elk37, Fie32, Mad37]. Native [Baj32b, Ban37b, Mad37, Riv37]. Natural [Baj35, Foe31, Haz33, Raw38, Smi37a, Sur35b]. Necessity [Rad38]. Need [Cla34, HT35, Lan35a]. nelsoni [Nee38]. Net [Wes32]. Nets [Ric32]. Newfoundland [Bel37a]. Newly [Wic32]. Nine [Boe38]. No [Bos37, Rob37]. Nocturnal [Wie34]. North [HD34, She37]. North-Eastern [HD34]. Northeast [vIW35]. Note [Smi35a, Wer34]. Notes [BH32, Hay32b, Lan35b, Lor32, NB33, Nee38, Sav35b, Smi37b, War31, Wie30, Wie31]. Nova [Smi35a, Smi38]. Nursery [Ait34]. Nutrition [Hea30, Hew37]. Nutritional [MTC^{+}30, MTC^{+}31, TM33, TM35]. Nutritive [CF34, WF35].

Objectives [Mot35]. Observations [Bel35a, BC38, Lan37b, MW34, NT34, Nur31, Ols32, Ric34, Ric37, Ric38, Smi37a, SA32, Sur33a, SO37, Whi30]. Ocean [BH32]. Oceanographic [Bor32]. Oceanographical [Hac37]. ocellatus [WF35]. Octomitus [Ric35]. off [Pri32]. Officers [Ano30e, Ano31e, Ano32e, Ano33e, Ano34d, Ano35d, Ano37d, Ano38d, Cle31]. Ohio [Ano33g, Ban37a, Ban33, Lan35b, Lan38, Roa33, Tra32, Tra33, Wic31, Wic32, Wic33b, WR37, Wic38]. Oil [Fie30, Gow35, WBF34, Wie35]. Oklahoma [Wic35]. One [Sur33b, Sur37]. Ontario [Ano30a, Ber30, Har30, Mac30a, Mac30b, McK30, Ric30, Sny32, Tes30]. Open [Bel31]. Operated [Coo37]. Operations [Kin34]. Organisms [ES34, SM31]. Origin [Mot34]. Osmerus [RB34, Van37b]. Other [ES34]. Otselic [Kin34, KR35]. Otter [Her32]. Our [Fie37, Gor33, Mar34, Mil32, Haw33]. Outlet [Ait34, CT30]. Oxygen [RW34, Smi34, Wie31, WM32, WF33, Wie34]. Oyster [Pry32, Rad33]. Oysters [Gow35].

Pacific [Bel33, Duf32, Pri32]. Panacea [Fie32]. Paper [Bru30]. Parasite [Tid34]. Parasites [Ban33, Ban37a, Par34, Ric37, War31, Wol35]. Parks [Mad37]. Parr [BPR32, Bel34b, Bel35a, Bel37a, BC38]. Passing [Har32]. Past [Hun38]. Pearl [Van37a]. Penaeus [WLA32]. Perch [Hil37]. Perches [Dea33]. Period [Whi34b]. Periods [Emb34, Van37a]. Permanganate [Pre34a, Wol35]. Phases [Haz32]. Physical [Roa33]. Physiography [Hac37]. Physiological [Wer34]. Pike [But37, Dea33, Hil37, Raw32, TH35]. Pike-Perch [Hil37]. Pike-Perches [Dea33]. Pikes [McN37]. Pimephales [WE32]. Pittsford [Lor32]. Places [Ano30f, Ano31f, Ano32f, Ano33f, Ano34e, Ano35e, Ano37e, Ano38e]. Plankton [Phi34, Ric32, Sur30b]. Planned [Lan35a]. Plant [Nur30, Ols32]. Planting [Foe34, RB34]. Plantings [SA32]. Plants [Cob33, RW34, Spi34]. Platyonichus [WF35]. Policy [Tra32]. Pollution [Ber30, Ell35, Gow35, Mac30a, McK30, Tra33, Wic30, WBF34]. Pomolobus [Ode34]. Pond [Ait34, Ald35, CT30, Lan31, Lan33, Lan37a, Mot35, Din32]. Pond-Fish [Lan33]. Pondfish [Lan32, Mee33, Mee34]. Ponds [Coo37, ES34, Fai34, Fis33, Hay32a, Lan34a, Lan35b, Mee35, Mot37, Smi32, Sur31a, SO37, Wie34]. Pool [Sur33a]. Pools [SA32]. Population [Baj32a, Joh34, She37, Smi35a, Smi35b]. Populations [Smi38, Tes33, VW38b]. Possible [Dym30]. Potassium [Pre34a, Wol35]. Power [Foe30]. Practical [CF34, Gow32b, Tra32, Wic33a]. Practices [FS35]. Prairie [Baj30, But30]. Precise [Gow32b]. Predator [Cle34, Riv37, Sal33]. Predators [Gre32]. Predatory [Smi35b]. Preliminary [BM35, Cro30, Dea33, Haz35, Raw32, Smi35a, Smi38, Van30]. Present [Fie30, Mac30b]. Preservation [Cun35a]. Preservatives [Lin30]. Presidents’ [Ano30f, Ano31f, Ano32f, Ano33f, Ano34e, Ano35e, Ano37e, Ano38e]. Prince [Whi30]. Principles [Dav30, Osb33, Sch37b, Vio31a]. Probable [Bir32]. Problem [Cle30, Cle34, Gre32, Hun30, Hun34b, Lan31, MGG^{+}34, WBF34]. Problems [Age33, Ald35, Ell37, Fie32, Hac37, Lan32, Lan33, Mac30a, Rus37]. Production [Cun35a, Dym30, Foe34, Hay33, Haz33, Lan31, Lan34b, Mee34, Sur33b, Sur37, Tar38]. Productive [Haz35, Rod30]. Productivity [Baj32a, Cun35b, Vio35]. Products [Don35]. Program [Bra35, Cah37, Lan38, Moo31]. Programs [Pri32]. Progress [Cul32, Gre35, HE30, Mar33b]. Prolonged [WM32]. promelas [WE32]. Propagating [ES34]. Propagation [Bro31, But37, Foe31, Hil37, How32, KR35, Mar33b, Mot32, Rad31, Shi31, Smi37a]. Proper [Van37c]. Prophylaxis [Atk32a]. Proposed [Bel32]. Protection [Mad37]. Province [Spi34, Mac30b, Sc.32, Wat32]. Provinces [Baj30, But30]. pseudaplites [Ban33, How31, Vio31b]. pseudoharengus [Ode34]. pulex [Smi37b]. punctatus [Mob31]. Pure [Wie31]. Purpose [Dav34b].

Qualitative [Age34d]. Quantitative [Age34d, HT34, Nee34, Sur30b, Sur30a, Sur33b]. Quebec [Spi34, Ric35, Ric37, Sc.32, Wat32]. Questionnaires [Van34]. Quick [Bir32]. Quick-Freezing [Bir32].

Raceways [Sav37]. Raf. [WE32]. Rafinesque [Smi35c]. Rainbow [Age34c, Atk32b, Gre32, Gre33, Mot33, Mot34, Mot38a, Mot38b, SA32, Sur33b, Sur37]. Raising [Hay32b, Pow31]. Rana [Vio31a]. Rapid [Pre34b]. Rapids [Bel37b]. Rate [Coo37, Dea33, Gre33, She37]. Rations [Foe32]. Reactions [Foe30, Jam30]. Reared [Age34d]. Rearing [Age33, CT30, Cob31, Cob32, Coo37, Hog33, Hog34, Kin37, Lan34a, Lan37a, Lin30, Lor30, Raw38, Sur33c]. Record [Ric35]. Recording [Bel32]. Reference [Age33, Lor33, RA34, Wie34, WBF34]. Region [Smi37a, Van37b, VW38b]. Regulation [Bel33, Mén38]. Regulations [Van34]. Relation [Cle34, Emb34, Gow32a, Gow35, Hil37, Hun35, Lin30, M’G32, Pri32, Wes32]. Relations [Mot34]. Relationship [RW34, Van37a]. Relationships [Ode34]. Relative [GS34]. Remarkable [Cun35b]. Remarks [VW38b]. Remedial [Cul30]. Remedies [Fie30]. Report [BM35, Cro30, Dea33, Ewe33, Raw32, Sav37, Van30]. Reported [McK30]. Requirements [JSR33, MTC^{+}30, MTC^{+}31, TM33, TM35]. Research [Bel38b, Cle32, Har30, Hub30, Sch37b, Van35, Wic31, Wic33c]. Reservoir [Moo31]. Resources [Cul30, Puj38]. Responds [Ear37]. Results [Atk32b, But30, Cob33, Coo35]. Retain [Mar35]. Return [Hun34a]. Returns [Wic33b, Wic38]. Ring [Han37]. River [McK30, She37, Smi37c, BC38, Bel38a, Cul30, Dou35, Fol32, Wic30]. Rivers [Bel37a]. Roccus [VW38a]. Rochester [Mar33b]. Role [Mee33, Mee34]. Roller [Smi35c]. Run [BK34]. Russian [Bor34]. rusticus [Lan35b, Lan37b].

S. [Whi30, Whi34b]. Sable [She37]. Sacandaga [Moo31]. salar [Fol32]. Salient [Age33]. Saline [Wer34]. Salmo [Cur34, Fol32, Nee38]. salmoides [Coo37, WGS31]. Salmon [Age34b, BPR32, BH32, Bel34a, Bel34b, Bel34c, BK34, Bel35a, Bel37a, BC38, Bel38a, Cle30, Deu35, Don35, DF38, Foe30, Foe31, Foe32, Hun34a, Hun35, Pri32, Whi34a]. Salmonidae [BM35, CT30]. salmonis [LE31, Ric35]. Salmonoid [Age33, Mot35]. Salmonoids [Age34d]. Salt [WBF34]. Salvelinus [Haz32, Ric38]. Sampling [Mot37]. Sand [WF35]. Sandy [Nur30]. sapidissima [VW38b]. sapidus [WF35]. Saskatchewan [Raw32]. Satisfactory [Sea33]. Scale [Bel34b, RB34]. Scales [Han37]. Schaffer [LE31]. Scientific [Cle31, Gor33, Pry32, Tra32, Tru32, Van35]. Scotia [Smi35a]. Scotian [Hac37, LH35, Smi38]. Screening [Bel31]. Sculpins [Kos37]. Sea [HP37, Hun34a]. Season [Dav31, Esc37]. Seasonal [Nee30]. Second [Esc37]. Securing [Gow32b]. Selective [HE30, Shi34]. Seneca [Ode34]. Sept [Ano35b, Ano37b]. September [Ano31a, Ano32g, Ano33g, Bel37b, Wes34, Wic35, Ano38b]. Service [Ano30f, Ano31f, Ano32f, Ano33f, Ano34e, Ano35e, Ano37e, Ano38e]. setiferus [WLA32]. Several [Wie31, WM32]. Sewage [RA34]. Shad [VW38b]. Shed [TH35]. Shelf [Hac37, LH35]. Shellfish [Bir32, Ken32]. Ship [Fie37]. Showing [Ano30d, Ano31d, Ano32d, Ano33d, Ano34c, Ano35c, Ano37c, Ano38c, Moo32]. Shrimp [Gow32b, Sea33, Smi33, Tit30, WLA32]. Significance [Fie37, GS34]. Single [Lor33]. Situation [Fra38]. Six [McL30]. Sixtieth [Ano30a]. Sixty [Ano31a]. Sixty-First [Ano31a]. Skim [Deu35]. Slough [Sur30b]. Small [ES34, Esc38, Jam30, RB34, Sur30a, Tes30, Vio31b]. Small-mouth [Vio31b]. Small-Mouthed [Jam30, Tes30]. Smallmouth [Ban33, KR35, Kin37, Raw38, Sny32]. Smelt [Sch37a, Van37b]. Social [Lan34a]. Society [Ano32g, Ano33g, Ano38h, Bel37b, Wes34, Wic35, Ano35b, Ano37b, Ano38b, O’M32, Ano30a, Ano31a]. Sockeye [Cle30, Foe30, Foe32]. Sodium [Sur31a]. Some [Bor34, Bro31, But30, Cat30, Coo35, Cro30, Dav30, Din32, DF38, Dym30, Ell37, EB35, Fai34, Fie32, Foe30, Gre33, Haz32, Hub33, M’G32, M’G34, MW34, Ols32, Osb33, Roa33, Sch35, Sch37b, SO37, Whi30, Whi34a, WR37, Ode32, Mee34]. South [Kin34, KR35, WLA32]. Southern [Vio31b]. Southwestern [Tar38]. Spawn [Age34c]. Spawning [Bel34a, Bel34c, Dav31, Gre32, HT34, Jam30, Mot33, Mot38b, NT34, Nur31, Sch35, Tes30, Wer34, Whi34b]. Special [Age33, Ano34f, Ano35f, Ano37f, Ano38f, Lor33, RA34, Wie34, WBF34]. Species [Boe38, Emb34, Hil37, Hub31, Ode32, Smi37a, Spi34, Wie31, WM32]. Specimens [Loc30]. Speckled [Cat30, Hea30, Pre35, Ric37, Ric30, Sav35a, Shi34, Tay35, Wer34]. Sponge [Rad38]. Sport [CC33, TV37a]. Spot [Age33]. Spotted [Ban33, How32]. Spread [HP37]. Spring [BK34]. Spring-Run [BK34]. Springs [Ano31a]. St. [Fol32]. Standard [Lan34b]. Standing [Ano34g, Ano35g, Ano37g, Ano38g]. Staple [Nee33]. Starfish [Ken32]. State [Coo37, Deu34, HE30, Hay32a, Hay32b, Bel31, Bel35b, Ear37, Gre35]. State-Operated [Coo37]. States [WLA32]. Statistics [CC33, Gow32b, HD33, Sco32, Vio35]. Status [Haw33, Mac30b]. Steelhead [NT34, Pau38, Taf34]. Stizostedion [Dea33, Hil37]. Stock [Ear30, Lor30]. Stocked [HJ38]. Stocking [Atk32b, Bel32, Pre35, Tra32]. Stone [Smi35c]. Stream [Ait35, Ban37a, Ban37b, Dav34b, Dav35, Gre35, HTE33, Lor33, Lor35, MGG^{+}34, Nee34, Sur33b, Sur37, Tar31, Tar37, WBF34]. Streams [Ban33, Ear30, Ell35, Loc32, Loc34, Nee33, Ric34, Tar38]. Striped [TV37b, VW38a]. Studies [Gow32a, HT34, HM33, How31, LS37, Nee30, Nee34, Pau38, Roa33, Sal33, WR37]. Study [Atk32b, Deu34, Elk37, Haz35, Sur30b, Sur33b]. Studying [Sur30a]. Submarine [Hac37]. Submerged [Sur31a]. Substances [Mee35]. Successes [Sur31b]. Successful [Haz33, Web31]. Sucker [Hun35]. Suckers [Atk31, Joh34]. Sudden [Pow38, Wie31]. Suggested [Cul30, Den34, Har32]. Suggestions [Bor34, Cat30, Hub33, Lan35a]. Suitable [Wic32]. Sulphate [Cat34, Smi35b]. Summary [Ban33, Bor32, DF38, Ewe33, WLA32, Wic33c, WT33]. Summer [HE37]. Sunlight [Ols32]. Supply [Cob31, RW34, Tay35]. Survey [Dwy32, FS35, Mil37]. Surveys [Van34]. Survive [WM32]. Sustained [Van37c].

Table [DF38]. Tagged [Sny32, Wic33b, Wic38]. Tagging [Ban37b, Cob33, Pri32]. Tags [Mar33a, Nes33]. Taken [Lor33]. Tapeworm [Lan37a]. Technical [Bel34b]. Teeth [TH35]. Temperature [Cla33, Emb34, Haz33, M’G32, Mar32, Mar30, Nur31]. Tennessee [Cah37]. Terms [Ano30f, Ano31f, Ano32f, Ano33f, Ano34e, Ano35e, Ano37e, Ano38e]. Test [Lor35]. Tests [CF34, DF38, GS34]. Texas [Wie34, WBF34]. Text [Bos37]. Their [Par34, Van37c, Bro31, M’G32, RW34, TH35]. Theories [Whi34a]. Theory [Din35]. there [Bos37]. Things [Din32]. Think [Din32]. Thirty [McL30]. Thirty-Six [McL30]. Throughout [Lor33]. Time [Mot38b]. Toronto [Ano30a, Bel30]. Trade [Fie37]. Transactions [Ano32g, Ano33g, Ano38h, Bel37b, Wes34, Wic35, Ano30a, Ano31a]. Transgression [LH35]. Trap [Wes32]. Trapnets [Van32]. Trawling [Her32]. Treated [RA34]. Treatment [Dom35]. Trichodinid [Ric38]. Troughs [KR35, Lin30]. Trout [Age34b, Age34c, Atk31, Atk32a, Atk32b, Bea37, Bra35, Cat30, Cob31, Cob33, Cur34, Dav31, Dav32, Dav34a, Det30, Deu34, Deu35, Din34, DF38, Dou35, Dym30, Emb34, Fis33, Fis34, Gre32, Gre33, HE30, Hay32a, Hay32b, Haz32, Haz33, HM33, Haz35, Hea30, Hew37, HJ38, Hun35, LE31, LS37, Lin30, Lor30, Lor32, Lor33, Lor34, Lor35, Mar35, MTC^{+}30, MTC^{+}31, MGG^{+}34, Mot32, Mot33, Mot34, Mot38a, Mot38b, Nee30, NT34, Pau38, Pow31, Pre35, Ric37, Ric30, Sav35a, She37, Shi34, SA32, Sur33b, Sur33c, Sur35a, Sur35b, Sur37, Tar31, Tar37, TL34, Tay35, TM33, TM35, VD38, Wal35, Wer34, Whi30, Whi34b, Wol37a, Wol37b, HD34]. Tulsa [Ald35, Wic35]. Type [Fis35, Lor33].

Ulcer [Fis34]. Undeveloped [Puj38]. United [WLA32]. Unproductive [Rod30]. upon [BM35, Cro30, Mar33a, Mar30, Riv37]. Upper [Cul30, Cul34, Wic30]. Use [Dav32, Deu35, Don35, NB33, Smi35b, Wol35, Pre34a, Wer34]. Used [Coo32, Ewe33, Lan31, Lin30]. Utah [Haz35]. Utility [Ric32]. Utilization [BCF37, Lan35a].

Valley [Cah37]. Value [CF34, Dav34b, Loc34, RA34, Sch35, Tar37, Tra32, Van34, WF35, Wic33a]. Varying [Wie30]. Vegetation [Sur31a, Wie34]. Vermiculatus [McN37]. Vermont [Lor33, Lor35]. versus [Smi37a]. Vitamin [JSR33]. vitreum [Hil37]. vs [Din35]. Vt. [Lor32].

Walleyed [But37]. Warm [Tay35]. Warner [Bow37]. Warner-Bower [Bow37]. Washings [Pau38]. Washington [Bel31]. Waskesiu [Raw32]. Wastage [Fie30]. Water [Bel34d, Ber30, Cob31, Ell35, Fie30, Haz33, Haz35, Lan35a, Mar33a, NB33, Rad33, Ric32, Rod30, Rus37, Smi32, Smi37b, Tay35, Tit30, Wes32, Wie31, WBF34, Wie35]. Water-Farming [Rad33]. Waterfowl [RA34]. Waters [Bel32, Bor34, Ell37, Gre33, Haz33, LH35, LS37, Moo31, Moo32, Ric30, Roa33, Sal33, Smi34, Tes30, Tes33, Vio35, Wic32, WR37]. Watershed [Dou35]. Wayne [Ban37a]. Weather [Hac34]. Weight [Mot37, Mot38b]. West [Bel37a]. Western [Boe38, Ewe33, Fis35, WT33]. which [Mar32]. White [Age33, Han37]. White-Spot [Age33]. Whitefish [Cun35a, HD34, Wic33a]. Whiting [BCF37]. Wildlife [LeC30, RA34]. Wilson [Ode34]. Winnipeg [Cun35a, McK30]. Winter [HE37, HT35, Lor33, Tay35]. Wintering [Hay32a]. Wisconsin [HD34].

Wood [Lin30]. Woods [Bor32]. Work [Ait35, Haw33, HTE33, Tru32]. World [Fie37]. Worm [LE31].

Year [Ano30d, Ano31d, Ano32d, Ano33d, Ano34c, Ano35c, Ano37c, Ano38c, Kin34, Lor33, McL30, Sur30b]. Years [Dom35]. Yellow [Hil37]. Yield [Elk37]. Yields [Van37c]. York [Gre35, Deu34, Din32, Kin34, Kos37, Mar33b, Moo38, Ode32, Ode34, Sto38]. Young [Coo37, Don35, DF38, Wie30].

References

Agersborg:1933:SPA

[Age33] H. P. Kjershow Agersborg. Salient problems in the artificial rearing of salmonoid fishes, with special reference to intestinal fungisitosis and the cause of white-spot disease. Transactions of the American Fisheries Society, 63(1):240–250, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Agersborg:1934:AA

[Age34a] H. P. K. Agersborg. Aquiculture and agriculture. Transactions of the American Fisheries Society, 64(1):266–269, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Agersborg:1934:IBD

[Age34b] H. P. K. Agersborg. An inexpensive balanced diet for trout and salmon. Transactions of the American Fisheries Society, 64(1):155– 162, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Agersborg:1934:WDR

[Age34c] H. P. K. Agersborg. When do the rainbow trout spawn? Transactions of the American Fisheries Society, 64(1):167–169, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Agersborg:1934:QQA

[Age34d] H. P. Kjerschow Agersborg. Qualitative and quantitative analyses of foods for salmonoids reared artificially. Transactions of the American Fisheries Society, 64(1):435–442, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Aitken:1934:INP

[Ait34] W. W. Aitken. Iowa nursery pond kettle and outlet. Transactions of the American Fisheries Society, 64(1):170–171, January 1934. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Aitken:1935:ISI

[Ait35] W. W. Aitken. Iowa stream improvement work. Transactions of the American Fisheries Society, 65(1):322–323, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Aitken:1937:FMU

[Ait37] W. W. Aitken. Fish movement under the ice in an Iowa lake. Transactions of the American Fisheries Society, 66(1):98–103, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Aldrich:1935:TMF

[Ald35] A. D. Aldrich. The tulsa municipal fish hatchery and problems of pond fish culture. Transactions of the American Fisheries Society, 65(1): 179–183, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1930:TAF

[Ano30a] Anonymous. Transactions of the American Fisheries Society: Sixtieth Annual Meeting at Toronto, Ontario: August 27, 28, and 29, 1930. Transactions of the American Fisheries Society, 60(1):9–27, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1930:CI

[Ano30b] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 60(1):327, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1930:CLA

[Ano30c] Anonymous. Constitution and by-laws: (as amended to date). Transactions of the American Fisheries Society, 60(1):328–331, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1930:LMS

[Ano30d] Anonymous. List of members, 1930–1931: (showing year of election to membership). Transactions of the American Fisheries Society, 60(1):

332–350, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1930:O

[Ano30e] Anonymous. Officers for 1930–1931. Transactions of the American Fisheries Society, 60(1):3, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1930:PTS

[Ano30f] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 60(1):4–5, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1931:TAF

[Ano31a] Anonymous. Transactions of the American Fisheries Society: Sixty-First Annual Meeting at Hot Springs, Arkansas: September 21, 22, and 23, 1931. Transactions of the American Fisheries Society, 61 (1):9–33, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1931:CI

[Ano31b] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 61(1):283, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1931:CLA

[Ano31c] Anonymous. Constitution and by-laws: (as amended to date). Transactions of the American Fisheries Society, 61(1):284–287, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1931:LMS

[Ano31d] Anonymous. List of members, 1931–32: (showing year of election to membership). Transactions of the American Fisheries Society, 61(1): 288–308, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1931:O

[Ano31e] Anonymous. Officers for 1930–1931. Transactions of the American Fisheries Society, 61(1):3–4, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1931:PTS

[Ano31f] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 61(1):5–6, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1932:AFF

[Ano32a] Anonymous. Abstracts of foreign fishery literature. Transactions of the American Fisheries Society, 62(1):391–398, January 1932. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1932:CI

[Ano32b] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 62(1):399, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1932:CLA

[Ano32c] Anonymous. Constitution and by-laws: (as amended to date). Transactions of the American Fisheries Society, 62(1):400–402, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1932:LMS

[Ano32d] Anonymous. List of members: (showing year of election to membership). Transactions of the American Fisheries Society, 62(1):403–420, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1932:O

[Ano32e] Anonymous. Officers for 1930–1931. Transactions of the American Fisheries Society, 62(1):3–4, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1932:PTS

[Ano32f] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 62(1):5–6, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1932:TAF

[Ano32g] Anonymous. Transactions of the American Fisheries Society at Baltimore, Maryland: September 21, 22 and 23, 1932. Transactions of the American Fisheries Society, 62(1):11–39, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1933:CI

[Ano33a] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 63(1):417, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1933:C

[Ano33b] Anonymous. Committees, 1933–1934. Transactions of the American Fisheries Society, 63(1):4, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1933:CLA

[Ano33c] Anonymous. Constitution and by-laws: (as amended to date). Transactions of the American Fisheries Society, 63(1):418–420, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1933:LMS

[Ano33d] Anonymous. List of members: (showing year of election to membership). Transactions of the American Fisheries Society, 63(1):421–439, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1933:O

[Ano33e] Anonymous. Officers for 1932–1933. Transactions of the American Fisheries Society, 63(1):3, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1933:PTS

[Ano33f] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 63(1):5–6, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1933:TAF

[Ano33g] Anonymous. Transactions of the American Fisheries Society at Columbus, Ohio, September 18, 19 and 20, 1933. Transactions of the American Fisheries Society, 63(1):11–49, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1934:CI

[Ano34a] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 64(1):471, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1934:CLA

[Ano34b] Anonymous. Constitution and by-laws: (as amended to date). Transactions of the American Fisheries Society, 64(1):472–474, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1934:LMS

[Ano34c] Anonymous. List of members: (showing year of election to membership). Transactions of the American Fisheries Society, 64(1):475–494, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1934:O

[Ano34d] Anonymous. Officers for 1933–1934. Transactions of the American Fisheries Society, 64(1):3, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1934:PTS

[Ano34e] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 64(1):6–7, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1934:SCb

[Ano34f] Anonymous. Special committees. Transactions of the American Fisheries Society, 64(1):5, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1934:SCa

[Ano34g] Anonymous. Standing committees, 1934–1935. Transactions of the American Fisheries Society, 64(1):4, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1935:CI

[Ano35a] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 65(1):405, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1935:CLA

[Ano35b] Anonymous. Constitution and by-laws of the American Fisheries Society: (as amended sept. 11, 1935). Transactions of the American Fisheries Society, 65(1):406–409, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1935:LMS

[Ano35c] Anonymous. List of members: (showing year of election to membership). Transactions of the American Fisheries Society, 65(1):410–429, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1935:O

[Ano35d] Anonymous. Officers for 1934–1935. Transactions of the American Fisheries Society, 65(1):3, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1935:PTS

[Ano35e] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 65(1):6–7, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1935:SCb

[Ano35f] Anonymous. Special committees. Transactions of the American Fisheries Society, 65(1):5, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1935:SCa

[Ano35g] Anonymous. Standing committees, 1935–1936. Transactions of the American Fisheries Society, 65(1):4, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1937:CI

[Ano37a] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 66(1):445, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1937:CLA

[Ano37b] Anonymous. Constitution and by-laws of the American Fisheries Society: (as amended sept. 4, 1936). Transactions of the American Fisheries Society, 66(1):446–449, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1937:LMS

[Ano37c] Anonymous. List of members: (showing year of election to membership). Transactions of the American Fisheries Society, 66(1):450–467, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1937:O

[Ano37d] Anonymous. Officers for 1935–1936. Transactions of the American Fisheries Society, 66(1):3, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1937:PTS

[Ano37e] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 66(1):6–7, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1937:SCb

[Ano37f] Anonymous. Special committees. Transactions of the American Fisheries Society, 66(1):5, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1937:SCa

[Ano37g] Anonymous. Standing committees, 1936–1937. Transactions of the American Fisheries Society, 66(1):4, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1938:CI

[Ano38a] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 67(1):289, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1938:CLA

[Ano38b] Anonymous. Constitution and by-laws of the American Fisheries Society: (as amended September 4, 1936). Transactions of the American Fisheries Society, 67(1):290–293, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1938:LMS

[Ano38c] Anonymous. List of members: (showing year of election to membership). Transactions of the American Fisheries Society, 67(1):294–311, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1938:O

[Ano38d] Anonymous. Officers for 1936–1937. Transactions of the American Fisheries Society, 67(1):3, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1938:PTS

[Ano38e] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 67(1):6–7, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1938:SCb

[Ano38f] Anonymous. Special committees. Transactions of the American Fisheries Society, 67(1):5, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1938:SCa

[Ano38g] Anonymous. Standing committees, 1937–1938. Transactions of the American Fisheries Society, 67(1):4, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1938:TAM

[Ano38h] Anonymous. Transactions of the 67th Annual Meeting of the American Fisheries Society: Mexico City, Mexico, August 23, 24 and 25, 1937. Transactions of the American Fisheries Society, 67(1):13–39, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atkinson:1931:DGT

[Atk31] N. J. Atkinson. The destruction of grey trout eggs by suckers and bullheads. Transactions of the American Fisheries Society, 61(1): 183–188, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atkinson:1932:BDT

[Atk32a] N. J. Atkinson. A bacterial disease of trout eggs and fry and its prophylaxis with acriflavine. Transactions of the American Fisheries Society, 62(1):152–155, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atkinson:1932:SCR

[Atk32b] N. J. Atkinson. A study of comparative results from stocking barren lakes with rainbow trout. Transactions of the American Fisheries Society, 62(1):197–201, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bajkov:1930:FIF

[Baj30] Alexander Bajkov. Fishing industry and fisheries investigations in the Prairie Provinces. Transactions of the American Fisheries Society, 60(1):215–237, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bajkov:1932:FPP

[Baj32a] A. D. Bajkov. Fish population and productivity of lakes. Transactions of the American Fisheries Society, 62(1):307–316, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bajkov:1932:NGF

[Baj32b] Alexander Bajkov. Native game fish in Manitoba. Transactions of the American Fisheries Society, 62(1):377–379, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bajkov:1935:HED

[Baj35] A. D. Bajkov. How to estimate the daily food consumption of fish under natural conditions. Transactions of the American Fisheries Society, 65(1):288–289, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bangham:1933:PSB

[Ban33] Ralph V. Bangham. Parasites of the spotted bass, Micropterus pseudaplites Hubbs, and summary of parasites of smallmouth and largemouth black bass from Ohio streams. Transactions of the American Fisheries Society, 63(1):220–228, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bangham:1937:PWC

[Ban37a] R. V. Bangham. Parasites of Wayne County (Ohio) stream fish. Transactions of the American Fisheries Society, 66(1):357–358, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bangham:1937:TNS

[Ban37b] R. V. Bangham. Tagging native stream fish. Transactions of the American Fisheries Society, 66(1):316–323, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1938:OSP

[BC38] David L. Belding and Mildred P. Clark. Observations on the salmon parr of the Margaree River. Transactions of the American Fisheries Society, 67(1):184–194, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bean:1937:CUA

[BCF37] D. A. Bean, J. A. Clague, and C. R. Fellers. Composition and utilization of the Atlantic whiting, Merluccius bilinearis. Transactions of the American Fisheries Society, 66(1):415–421, January 1937. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beach:1937:DTF

[Bea37] U. Sidney Beach. The destruction of trout by fish ducks. Transactions of the American Fisheries Society, 66(1):338–342, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bell:1930:TAA

[Bel30] G. S. Bell. Toronto Anglers’ Association. Transactions of the American Fisheries Society, 60(1):238–239, January 1930. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bell:1931:SOD

[Bel31] Milo C. Bell. Screening of open diversions in the State of Washington. Transactions of the American Fisheries Society, 61(1):270–279, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1932:PMR

[Bel32] David L. Belding. A proposed method of recording fish growth and grading waters for stocking. Transactions of the American Fisheries Society, 62(1):256–257, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bell:1933:IRP

[Bel33] Frank T. Bell. The investigation and regulation of the Pacific halibut fishery. Transactions of the American Fisheries Society, 63(1):85–90, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1934:CHM

[Bel34a] David L. Belding. The cause of the high mortality in the Atlantic salmon after spawning. Transactions of the American Fisheries Society, 64(1):219–224, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1934:ITM

[Bel34b] David L. Belding. Improved technical methods for determining the annual growth of salmon parr by scale measurements. Transactions of the American Fisheries Society, 64(1):103–106, January 1934. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1934:SHA

[Bel34c] David L. Belding. The spawning habits of the Atlantic salmon. Transactions of the American Fisheries Society, 64(1):211–218, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bell:1934:WC

[Bel34d] Frank T. Bell. Water conservation. Transactions of the American Fisheries Society, 64(1):54–56, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1935:OGA

[Bel35a] David L. Belding. Observations on the growth of Atlantic salmon parr. Transactions of the American Fisheries Society, 65(1):157–160, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bell:1935:FSC

[Bel35b] Frank T. Bell. Federal and State cooperation in fishery investigations. Transactions of the American Fisheries Society, 65(1):246–252, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1937:ASP

[Bel37a] David L. Belding. Atlantic salmon parr of the West Coast rivers of Newfoundland. Transactions of the American Fisheries Society, 66(1): 211–224, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bell:1937:TAM

[Bel37b] Frank T. Bell. Transactions of the 66th Annual Meeting of the American Fisheries Society: Grand Rapids, Michigan, September 2, 3 and 4, 1936. Transactions of the American Fisheries Society, 66(1):13–48, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1938:SMR

[Bel38a] David L. Belding. The salmon of the Moisie River. Transactions of the American Fisheries Society, 67(1):195–206, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bell:1938:BIC

[Bel38b] Frank T. Bell. The benefits of international cooperation in fishery research. Transactions of the American Fisheries Society, 67(1):147– 154, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berry:1930:WPO

[Ber30] A. E. Berry. Water pollution in Ontario. Transactions of the American Fisheries Society, 60(1):306–310, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berriozabal:1938:MFH

[Ber38] Felipe B. Berriozabal. The management of fish hatcheries in Mexico. Transactions of the American Fisheries Society, 67(1):94–95, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1932:NOF

[BH32] David L. Belding and Doris M. Hyde. Notes on the ocean feeding grounds of the Atlantic salmon. Transactions of the American Fisheries Society, 62(1):304–306, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Birdseye:1932:PIQ

[Bir32] Clarence Birdseye. Probable influence of quick-freezing on the shellfish industries. Transactions of the American Fisheries Society, 62(1):80– 83, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1934:SRF

[BK34] David L. Belding and J. Arthur Kitson. Spring-run and fall-run Atlantic salmon. Transactions of the American Fisheries Society, 64(1): 225–230, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1935:PRU

[BM35] David L. Belding and Beulah Merrill. A preliminary report upon a hatchery disease of the Salmonidae. Transactions of the American Fisheries Society, 65(1):76–84, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boesel:1938:FNS

[Boe38] M. W. Boesel. The food of nine species of fish from the western end of Lake Erie. Transactions of the American Fisheries Society, 67(1): 215–223, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Borodin:1932:SEA

[Bor32] N. A. Borodin. Summary of experiments on the anabiosis of fishes made in 1931 and 1932 at Woods Hole Oceanographic Institution. Transactions of the American Fisheries Society, 62(1):340–341, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Borodin:1934:SSC

[Bor34] N. A. Borodin. Some suggestions concerning the introduction of some Russian fishes in Canadian waters. Transactions of the American Fisheries Society, 64(1):368–369, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bost:1937:WTN

[Bos37] J. E. Bost. Why is there no text book on fish culture? Transactions of the American Fisheries Society, 66(1):298–300, January 1937. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bower:1937:WBF

[Bow37] E. A. Bower. The Warner-Bower fish conveyor. Transactions of the American Fisheries Society, 66(1):411–414, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1932:EGS

[BPR32] David L. Belding, Mildred J. Pender, and James A. Rodd. The early growth of salmon parr in Canadian hatcheries. Transactions of the American Fisheries Society, 62(1):211–223, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Branion:1930:MLB

[Bra30] Hugh Branion. The marketing of ling (burbot). Transactions of the American Fisheries Society, 60(1):199–203, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bradt:1935:MBT

[Bra35] Glenn W. Bradt. Michigan’s beaver-trout management program. Transactions of the American Fisheries Society, 65(1):253–257, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1931:BAS

[Bro31] Dell Brown. The basses of Arkansas and some experiments in their propagation. Transactions of the American Fisheries Society, 61(1): 83–85, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bruce:1930:PF

[Bru30] C. Bruce. Paper on fishways. Transactions of the American Fisheries Society, 60(1):240–243, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burr:1931:EMG

[Bur31] J. G. Burr. Electricity as a means of garfish and carp control. Transactions of the American Fisheries Society, 61(1):174–182, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Butler:1930:FCP

[But30] G. E. Butler. Fish culture in the Prairie Provinces, and some of its results. Transactions of the American Fisheries Society, 60(1): 119–120, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Butler:1937:APW

[But37] George E. Butler. Artificial propagation of walleyed pike. Transactions of the American Fisheries Society, 66(1):277–278, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cahn:1937:FPT

[Cah37] Alvin R. Cahn. The fisheries program in the Tennessee Valley. Transactions of the American Fisheries Society, 66(1):398–402, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Catt:1930:SSB

[Cat30] James Catt. Some suggestions for breeding speckled trout. Transactions of the American Fisheries Society, 60(1):114–115, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Catt:1934:CSE

[Cat34] James Catt. Copper sulphate in the elimination of coarse fish. Transactions of the American Fisheries Society, 64(1):276–280, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:1933:MCS

[CC33] G. H. Clark and Richard Croker. A method of collecting statistics of marine sport catches in California. Transactions of the American Fisheries Society, 63(1):332–337, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cleveland:1934:CLP

[CF34] M. M. Cleveland and C. R. Fellers. Comparison of laboratory and practical tests for determining the nutritive value of fish meals. Transactions of the American Fisheries Society, 64(1):293–303, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clausen:1933:FMU

[Cla33] Ralph G. Clausen. Fish metabolism under increasing temperature. Transactions of the American Fisheries Society, 63(1):215–219, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:1934:NMA

[Cla34] G. H. Clark. The need for a measure of the Angler’s catch. Transactions of the American Fisheries Society, 64(1):49–53, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clemens:1930:PCS

[Cle30] W. A. Clemens. The problem of the conservation of the sockeye salmon in British Columbia. Transactions of the American Fisheries Society, 60(1):262–268, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clemens:1931:CSI

[Cle31] W. A. Clemens. A course of scientific instruction for fish cultural officers. Transactions of the American Fisheries Society, 61(1):164– 170, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clemens:1932:ARF

[Cle32] W. A. Clemens. The aim of research in fish culture in Canada. Transactions of the American Fisheries Society, 62(1):261–266, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clemens:1934:PCF

[Cle34] W. A. Clemens. The predator and coarse fish problem in relation to fish culture. Transactions of the American Fisheries Society, 64(1): 318–322, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cobb:1931:MML

[Cob31] Eben W. Cobb. Making the most of a limited water supply in trout rearing. Transactions of the American Fisheries Society, 61(1):39–42, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cobb:1932:IEL

[Cob32] Eben W. Cobb. Improved equipment for lobster rearing. Transactions of the American Fisheries Society, 62(1):84–87, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cobb:1933:RTT

[Cob33] Eben W. Cobb. Results of trout tagging to determine migrations and results from plants made. Transactions of the American Fisheries Society, 63(1):308–318, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cook:1932:MDU

[Coo32] A. B. Cook, Jr. Mechanical devices used in Michigan fish culture. Transactions of the American Fisheries Society, 62(1):179–184, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooper:1935:SRF

[Coo35] Gerald P. Cooper. Some results of forage fish investigations in Michigan. Transactions of the American Fisheries Society, 65(1):132–142, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooper:1937:FHR

[Coo37] Gerald P. Cooper. Food habits, rate of growth and cannibalism of young largemouth bass (Aplites salmoides) in state-operated rearing ponds in Michigan during 1935. Transactions of the American Fisheries Society, 66(1):242–266, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Crowell:1930:PRU

[Cro30] Mary Crowell. A preliminary report upon some irradiation experiments. Transactions of the American Fisheries Society, 60(1):167– 171, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Croker:1937:HCM

[Cro37] Richard S. Croker. How California is measuring the angler’s catch. Transactions of the American Fisheries Society, 66(1):301–305, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cobb:1930:CPC

[CT30] Eben W. Cobb and J. W. Titcomb. A circular pond with central outlet for rearing fry and fingerlings of the Salmonidae. Transactions of the American Fisheries Society, 60(1):121–123, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Culler:1930:DAR

[Cul30] C. F. Culler. Depletion of the aquatic resources of the Upper Mississippi River and suggested remedial measures. Transactions of the American Fisheries Society, 60(1):279–283, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Culler:1932:PFC

[Cul32] C. F. Culler. Progress in fish culture. Transactions of the American Fisheries Society, 62(1):114–118, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Culler:1934:FUM

[Cul34] C. F. Culler. The future of the Upper Mississippi fisheries. Transactions of the American Fisheries Society, 64(1):328–331, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cunningham:1935:PWP

[Cun35a] A. G. Cunningham. The preservation of whitefish production in Lake Winnipeg. Transactions of the American Fisheries Society, 65(1): 272–274, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cunningham:1935:RPL

[Cun35b] A. G. Cunningham. Remarkable productivity of Lake Dauphin. Transactions of the American Fisheries Society, 65(1):275–276, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Curtis:1934:GTC

[Cur34] Brian Curtis. The golden trout of Cottonwood Lakes (Salmo aguabonita Jordan). Transactions of the American Fisheries Society, 64(1): 259–265, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1930:SPB

[Dav30] H. S. Davis. Some principles of bass culture. Transactions of the American Fisheries Society, 60(1):48–52, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1931:IHS

[Dav31] H. S. Davis. The influence of heredity on the spawning season of trout. Transactions of the American Fisheries Society, 61(1):43–47, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1932:UDF

[Dav32] H. S. Davis. The use of dry foods in the diet of trout. Transactions of the American Fisheries Society, 62(1):189–196, January 1932. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1934:GHT

[Dav34a] H. S. Davis. Growth and heredity in trout. Transactions of the American Fisheries Society, 64(1):197–202, January 1934. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1934:PVS

[Dav34b] H. S. Davis. The purpose and value of stream improvement. Transactions of the American Fisheries Society, 64(1):63–67, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1935:SMN

[Dav35] H. S. Davis. Stream management in the national forests. Transactions of the American Fisheries Society, 65(1):234–239, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deason:1933:PRG

[Dea33] Hilary J. Deason. Preliminary report on the growth rate, dominance, and maturity of the pike-perches (Stizostedion) of Lake Erie. Transactions of the American Fisheries Society, 63(1):348–360, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Denmead:1931:FBB

[Den31] Talbott Denmead. The Federal Black Bass Law. Transactions of the American Fisheries Society, 61(1):258–261, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Denmead:1934:SBB

[Den34] Talbott Denmead. Suggested black bass legislation, 1935. Transactions of the American Fisheries Society, 64(1):97–102, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Detwiler:1930:FEB

[Det30] J. D. Detwiler. Feeding experiments with brook trout fingerlings. Transactions of the American Fisheries Society, 60(1):146–157, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deuel:1934:SFC

[Deu34] Charles R. Deuel. A study of the food cost of trout in New York state hatcheries. Transactions of the American Fisheries Society, 64(1): 172–175, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deuel:1935:UFM

[Deu35] Charles R. Deuel. The use of fish meal, cottonseed meal, meat meal, salmon egg meal and dried skim milk in trout diets. Transactions of the American Fisheries Society, 65(1):161–164, January 1935. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Donaldson:1938:STS

[DF38] Lauren R. Donaldson and Fred J. Foster. A summary table of some experimental tests in feeding young salmon and trout. Transactions of the American Fisheries Society, 67(1):262–270, January 1938. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dinsmore:1932:STW

[Din32] A. H. Dinsmore. Some things we think we have learned at York Pond. Transactions of the American Fisheries Society, 62(1):141–143, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dinsmore:1934:EHG

[Din34] A. H. Dinsmore. Effect of heredity on the growth of brook trout. Transactions of the American Fisheries Society, 64(1):203–204, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dinsmore:1935:FVT

[Din35] A. H. Dinsmore. Fact vs. theory. Transactions of the American Fisheries Society, 65(1):204–206, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Domogalla:1935:EYC

[Dom35] Bernhard Domogalla. Eleven years of chemical treatment of the Madison Lakes: — its effect on fish and fish foods. Transactions of the American Fisheries Society, 65(1):115–121, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Donaldson:1935:USP

[Don35] Lauren R. Donaldson. The use of salmon by-products as food for young king salmon. Transactions of the American Fisheries Society, 65(1):165–171, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Doudoroff:1935:FTL

[Dou35] Peter Doudoroff. Food of trout from lakes in the Klamath River watershed. Transactions of the American Fisheries Society, 65(1):129–131, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Duff:1932:FPC

[Duf32] D. C. B. Duff. Furunculosis on the Pacific Coast. Transactions of the American Fisheries Society, 62(1):249–255, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dwyer:1932:HBF

[Dwy32] J. N. Dwyer. Hydrographic bottle for fisheries survey. Transactions of the American Fisheries Society, 62(1):342–343, January 1932. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dymond:1930:PCF

[Dym30] J. R. Dymond. A possible critical factor affecting the production of trout in some British Columbia lakes. Transactions of the American Fisheries Society, 60(1):247–249, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Earle:1930:WMD

[Ear30] Swepson Earle. What Maryland is doing to stock her streams. Transactions of the American Fisheries Society, 60(1):260–261, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Earle:1932:FCB

[Ear32] Swepson Earle. The fisheries of Chesapeake Bay. Transactions of the American Fisheries Society, 62(1):43–49, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Earle:1937:FIR

[Ear37] Swepson Earle. Fish increase responds to Government and State cooperation. Transactions of the American Fisheries Society, 66(1):

392–394, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ewers:1935:FSB

[EB35] Lela A. Ewers and M. W. Boesel. The food of some Buckeye Lake fishes. Transactions of the American Fisheries Society, 65(1):57–70, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Elkins:1937:FYS

[Elk37] Winston A. Elkins. A fish yield study for certain lakes in the Chequamegon National Forest. Transactions of the American Fisheries Society, 66(1):306–312, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ellis:1935:MPF

[Ell35] M. M. Ellis. Measuring pollution in fresh water streams. Transactions of the American Fisheries Society, 65(1):240–245, January 1935. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ellis:1937:SFP

[Ell37] M. M. Ellis. Some fishery problems in impounded waters. Transactions of the American Fisheries Society, 66(1):63–75, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Embody:1934:RTI

[Emb34] G. C. Embody. Relation of temperature to the incubation periods of eggs of four species of trout. Transactions of the American Fisheries Society, 64(1):281–292, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Embody:1934:PDO

[ES34] G. C. Embody and W. O. Sadler. Propagating Daphnia and other forage organisms intensively in small ponds. Transactions of the American Fisheries Society, 64(1):205–210, January 1934. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eschmeyer:1935:AGF

[Esc35] R. W. Eschmeyer. Analysis of the game-fish catch in a Michigan lake. Transactions of the American Fisheries Society, 65(1):207–223, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eschmeyer:1937:SSC

[Esc37] R. W. Eschmeyer. A second season of creel census on Fife Lake. Transactions of the American Fisheries Society, 66(1):324–334, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eschmeyer:1938:EMG

[Esc38] R. William Eschmeyer. Experimental management of a group of small Michigan lakes. Transactions of the American Fisheries Society, 67(1): 120–129, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ewers:1933:SRC

[Ewe33] Lela A. Ewers. Summary report of crustacea used as food by the fishes of the western end of Lake Erie. Transactions of the American Fisheries Society, 63(1):379–390, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Faigenbaum:1934:SCC

[Fai34] Harold M. Faigenbaum. Some chemical characteristics of Adirondack lakes and ponds. Transactions of the American Fisheries Society, 64(1):189–196, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Field:1930:PFW

[Fie30] George W. Field. Present and future wastage of oil and water: Effects and remedies. Transactions of the American Fisheries Society, 60(1): 284–285, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Field:1932:BPS

[Fie32] George W. Field. Biology as the panacea for some national problems. Transactions of the American Fisheries Society, 62(1):369–376, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fielder:1937:FSS

[Fie37] R. H. Fielder. The factory ship — its significance to our world trade and commerce. Transactions of the American Fisheries Society, 66(1): 427–441, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fish:1933:CDT

[Fis33] Frederic F. Fish. The chemical disinfection of trout ponds. Transactions of the American Fisheries Society, 63(1):158–163, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fish:1934:UDT

[Fis34] Frederic F. Fish. Ulcer disease of trout. Transactions of the American Fisheries Society, 64(1):252–258, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fish:1935:WTB

[Fis35] Frederic F. Fish. A western type of bacterial gill disease. Transactions of the American Fisheries Society, 65(1):85–87, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Foerster:1930:RSS

[Foe30] R. E. Foerster. The reactions of some sockeye salmon to a power dam. Transactions of the American Fisheries Society, 60(1):244–246, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Foerster:1931:CNA

[Foe31] R. E. Foerster. A comparison of the natural and artificial propagation of salmon. Transactions of the American Fisheries Society, 61(1): 121–130, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Foerster:1932:AFR

[Foe32] R. E. Foerster. Artificial food rations for sockeye salmon. Transactions of the American Fisheries Society, 62(1):228–238, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Foerster:1934:FPE

[Foe34] R. E. Foerster. Fry production from eyed-egg planting. Transactions of the American Fisheries Society, 64(1):379–381, January 1934. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Follett:1932:SSS

[Fol32] Richard Follett. Salmo salar of the St. Lawrence River. Transactions of the American Fisheries Society, 62(1):366–368, January 1932. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fragoso:1938:SFI

[Fra38] Francisco Navarro Fragoso. The situation of the fishing industry in Mexico. Transactions of the American Fisheries Society, 67(1):256– 261, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fiedler:1935:SFH

[FS35] R. H. Fiedler and V. J. Samson. Survey of fish hatchery foods and feeding practices. Transactions of the American Fisheries Society, 65(1):376–402, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gordon:1933:SMO

[Gor33] Seth Gordon. Scientific management — our future fisheries job. Transactions of the American Fisheries Society, 63(1):73–79, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gowanloch:1932:ICH

[Gow32a] James Nelson Gowanloch. The importance and conduct of hydrographic studies in Louisiana in relation to commercial fisheries. Transactions of the American Fisheries Society, 62(1):336–339, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gowanloch:1932:PIS

[Gow32b] James W. Gowanloch. Practical importance of securing more precise statistics of the shrimp fishery. Transactions of the American Fisheries Society, 62(1):111–113, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gowanloch:1935:POR

[Gow35] James Nelson Gowanloch. Pollution by oil in relation to oysters. Transactions of the American Fisheries Society, 65(1):293–296, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Greeley:1932:SHB

[Gre32] John R. Greeley. The spawning habits of brook, brown and rainbow trout, and the problem of egg predators. Transactions of the American Fisheries Society, 62(1):239–248, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Greeley:1933:GRR

[Gre33] John R. Greeley. The growth rate of rainbow trout from some Michigan waters. Transactions of the American Fisheries Society, 63(1): 361–378, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Greeley:1935:PSI

[Gre35] John R. Greeley. Progress of stream improvement in New York State. Transactions of the American Fisheries Society, 65(1):316–321, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Griffiths:1934:SBC

[GS34] Francis P. Griffiths and Maurice E. Stansby. Significance of the bacterial count and chemical tests in determining the relative freshness of haddock. Transactions of the American Fisheries Society, 64(1): 401–407, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hachey:1934:WMC

[Hac34] H. B. Hachey. The weather man and coastal fisheries. Transactions of the American Fisheries Society, 64(1):382–389, January 1934. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hachey:1937:SPO

[Hac37] H. B. Hachey. The submarine physiography and oceanographical problems of the Scotian Shelf. Transactions of the American Fisheries Society, 66(1):237–241, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:1937:DAR

[Han37] Donald F. Hansen. The date of annual ring formation in the scales of the white crappie. Transactions of the American Fisheries Society, 66(1):227–236, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harkness:1930:OFR

[Har30] William J. K. Harkness. Ontario fisheries research laboratory. Transactions of the American Fisheries Society, 60(1):178–181, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harrison:1932:SMP

[Har32] C. W. Harrison. Suggested method of passing fish over high dams. Transactions of the American Fisheries Society, 62(1):386–388, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hawes:1933:SOB

[Haw33] Harry B. Hawes. The status of our black bass work. Transactions of the American Fisheries Society, 63(1):98–102, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayford:1932:EBC

[Hay32a] Charles O. Hayford. Economic bass culture in trout wintering ponds at the Hackettstown State Fish Hatchery. Transactions of the American Fisheries Society, 62(1):167–170, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayford:1932:FCN

[Hay32b] Charles O. Hayford. Fish cultural notes on trout raising in the New Jersey State Fish Hatchery. Transactions of the American Fisheries Society, 62(1):126–132, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayford:1933:PLB

[Hay33] Charles O. Hayford. Production of largemouth bass. Transactions of the American Fisheries Society, 63(1):164–166, January 1933. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hazzard:1932:SPL

[Haz32] A. S. Hazzard. Some phases of the life history of the Eastern brook trout, Salvelinus fontinalis Mitchell. Transactions of the American Fisheries Society, 62(1):344–350, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hazzard:1933:LWT

[Haz33] A. S. Hazzard. Low water temperature, a limiting factor in the successful production of trout in natural waters. Transactions of the American Fisheries Society, 63(1):204–207, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hazzard:1935:PSE

[Haz35] A. S. Hazzard. A preliminary study of an exceptionally productive trout water, Fish Lake, Utah. Transactions of the American Fisheries Society, 65(1):122–128, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hile:1933:MIS

[HD33] Ralph Hile and William R. Duden. Methods for the investigation of the statistics of the commercial fisheries of the Great Lakes. Transactions of the American Fisheries Society, 63(1):292–305, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hile:1934:GWC

[HD34] Ralph Hile and Hilary J. Deason. Growth of the whitefish, Coregonus clupeaformis (Mitchill), in Trout Lake, North-Eastern Highlands, Wisconsin. Transactions of the American Fisheries Society, 64(1):231–237, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayford:1930:FPS

[HE30] Charles O. Hayford and George C. Embody. Further progress in the selective breeding of brook trout at the New Jersey State Hatchery. Transactions of the American Fisheries Society, 60(1):109–113, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hazzard:1937:CSW

[HE37] A. S. Hazzard and R. W. Eschmeyer. A comparison of summer and winter fishing in Michigan lakes. Transactions of the American Fisheries Society, 66(1):87–97, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heard:1930:INS

[Hea30] R. D. H. Heard. Investigations on the nutrition of speckled trout. Transactions of the American Fisheries Society, 60(1):140–145, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Herrington:1932:CIF

[Her32] William C. Herrington. Conservation of immature fish in otter trawling. Transactions of the American Fisheries Society, 62(1):57–63, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hewitt:1937:IFH

[Hew37] Edward R. Hewitt. Identification of factor “H” in the nutrition of trout. Transactions of the American Fisheries Society, 66(1):291–297, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hile:1937:IAY

[Hil37] Ralph Hile. The increase in the abundance of the yellow pike-perch, Stizostedion vitreum (Mitchill), in Lakes Huron and Michigan, in relation to the artificial propagation of the species. Transactions of the American Fisheries Society, 66(1):143–159, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoover:1938:MDS

[HJ38] Earl E. Hoover and M. S. Johnson. Migration and depletion of stocked brook trout. Transactions of the American Fisheries Society, 67(1): 224–227, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hazzard:1933:SFC

[HM33] A. S. Hazzard and M. J. Madsen. Studies of the food of the cutthroat trout. Transactions of the American Fisheries Society, 63(1):198–203, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hogan:1933:ECF

[Hog33] Joe Hogan. Experiments with commercial fertilizers in rearing largemouth black bass fingerlings. Transactions of the American Fisheries Society, 63(1):110–119, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hogan:1934:RLB

[Hog34] Joe Hogan. Rearing largemouth black bass at lonoke. Transactions of the American Fisheries Society, 64(1):127–131, January 1934. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Howland:1931:SKB

[How31] Joe W. Howland. Studies on the Kentucky black bass: (Micropterus pseudaplites Hubbs). Transactions of the American Fisheries Society, 61(1):89–94, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Howland:1932:EPS

[How32] Joe W. Howland. Experiments in the propagation of spotted black bass. Transactions of the American Fisheries Society, 62(1):185–188, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hubbs:1937:SSL

[HP37] Carl L. Hubbs and T. E. B. Pope. The spread of the sea lamprey through the Great Lakes. Transactions of the American Fisheries Society, 66(1):172–176, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hart:1934:QSH

[HT34] John Lawson Hart and Albert L. Tester. Quantitative studies on herring spawning. Transactions of the American Fisheries Society, 64(1):307–312, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hubbs:1935:NIF

[HT35] Carl L. Hubbs and Milton B. Trautman. The need for investigating fish conditions in winter. Transactions of the American Fisheries Society, 65(1):51–56, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hubbs:1933:CCC

[HTE33] Carl L. Hubbs, Clarence M. Tarzwell, and R. W. Eschmeyer. C.C.C. stream improvement work in Michigan. Transactions of the American Fisheries Society, 63(1):404–416, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hubbs:1930:FRM

[Hub30] Carl L. Hubbs. Fishery research in Michigan. Transactions of the American Fisheries Society, 60(1):182–186, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hubbs:1931:SBB

[Hub31] Carl L. Hubbs. The species of black bass. Transactions of the American Fisheries Society, 61(1):86–88, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hubbs:1933:SES

[Hub33] Carl L. Hubbs. Some experiences and suggestions on forage fish culture. Transactions of the American Fisheries Society, 63(1):53–63, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hubbs:1937:FML

[Hub37] Carl L. Hubbs. Fish management: Looking forward. Transactions of the American Fisheries Society, 66(1):51–55, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huntsman:1930:AFP

[Hun30] A. G. Huntsman. Attack on the fishery problem. Transactions of the American Fisheries Society, 60(1):172–177, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huntsman:1934:FIR

[Hun34a] A. G. Huntsman. Factors influencing return of salmon from the sea. Transactions of the American Fisheries Society, 64(1):351–355, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huntsman:1934:PCA

[Hun34b] A. G. Huntsman. The problem of control in aquiculture. Transactions of the American Fisheries Society, 64(1):364–367, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huntsman:1935:SCC

[Hun35] A. G. Huntsman. The sucker (Catostomus commersonii) in relation to salmon and trout. Transactions of the American Fisheries Society, 65(1):152–156, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huntsman:1938:FCP

[Hun38] Archibald G. Huntsman. Fish culture — past and future. Transactions of the American Fisheries Society, 67(1):87–93, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

James:1930:SRS

[Jam30] M. C. James. Spawning reactions of small-mouthed bass. Transactions of the American Fisheries Society, 60(1):62–63, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

James:1934:EDF

[Jam34] M. C. James. Effect of 1934 drought on fish life. Transactions of the American Fisheries Society, 64(1):57–62, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnston:1934:CBS

[Joh34] Herbert Johnston. Can blood suckers affect the fish population of a lake? Transactions of the American Fisheries Society, 64(1):363, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jewell:1933:VRG

[JSR33] Minna E. Jewell, Edward Schneberger, and Jean Alexander Ross. The vitamin requirements of goldfish and channel cat. Transactions of the American Fisheries Society, 63(1):338–347, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kenney:1932:CSM

[Ken32] Raymond J. Kenney. Combating the starfish menace on Massachusetts shellfish grounds. Transactions of the American Fisheries Society, 62 (1):76–79, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kingsbury:1934:NBH

[Kin34] Oliver R. Kingsbury. The new bass hatchery at South Otselic, New York, and its first year’s operations. Transactions of the American Fisheries Society, 64(1):132–142, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kingsbury:1937:FER

[Kin37] Oliver R. Kingsbury. Foes encountered in the rearing of smallmouth bass. Transactions of the American Fisheries Society, 66(1):267–274, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Koster:1937:FSC

[Kos37] Wm J. Koster. The food of sculpins (Cottidae) in Central New York. Transactions of the American Fisheries Society, 66(1):374–382, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kingsbury:1935:PSB

[KR35] O. R. Kingsbury and Wm F. Royce. Propagation of smallmouth bass in troughs at South Otselic Bass Hatchery. Transactions of the American Fisheries Society, 65(1):309–312, January 1935. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Langlois:1931:PEM

[Lan31] T. H. Langlois. The problem of efficient management of hatcheries used for the production of pond fishes. Transactions of the American Fisheries Society, 61(1):106–115, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Langlois:1932:PPC

[Lan32] T. H. Langlois. Problems of pondfish culture. Transactions of the American Fisheries Society, 62(1):156–166, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Langlois:1933:MPP

[Lan33] T. H. Langlois. More problems of pond-fish culture. Transactions of the American Fisheries Society, 63(1):391–403, January 1933. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Langlois:1934:SBB

[Lan34a] T. H. Langlois. The social behavior of bass in rearing ponds. Transactions of the American Fisheries Society, 64(1):146–150, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Langlois:1934:SMC

[Lan34b] T. H. Langlois. Standard methods of computing bass production. Transactions of the American Fisheries Society, 64(1):163–166, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Langlois:1935:NPW

[Lan35a] T. H. Langlois. The need for planned water utilization with aquicultural suggestions. Transactions of the American Fisheries Society, 65(1):284–287, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Langlois:1935:NHC

[Lan35b] T. H. Langlois. Notes on the habits of the crayfish, Cambarus rusticus Girard, in fish ponds in Ohio. Transactions of the American Fisheries Society, 65(1):189–193, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Langlois:1937:BTI

[Lan37a] T. H. Langlois. Bass tapeworm infection in a rearing pond. Transactions of the American Fisheries Society, 66(1):364–366, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Langlois:1937:FOH

[Lan37b] T. H. Langlois. Further observations on habits of the crayfish, Cambarus rusticus Girard. Transactions of the American Fisheries Society, 66(1):275–276, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Langlois:1938:OFM

[Lan38] T. H. Langlois. The Ohio fisheries management program. Transactions of the American Fisheries Society, 67(1):114–119, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Laird:1931:CTG

[LE31] James Laird and G. C. Embody. Controlling the trout gill worm (Discocotyle salmonis, Schaffer). Transactions of the American Fisheries Society, 61(1):189–192, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

LeCompte:1930:EDW

[LeC30] E. Lee LeCompte. Effect of drought on wildlife. Transactions of the American Fisheries Society, 60(1):251–252, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leim:1935:TMW

[LH35] A. H. Leim and H. B. Hachey. A transgression of marginal waters over the Scotian Shelf. Transactions of the American Fisheries Society, 65(1):279–283, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lincoln:1930:HGT

[Lin30] Guy Lincoln. Health and growth of trout with relation to wood preservatives used in rearing troughs. Transactions of the American Fisheries Society, 60(1):124–126, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Locke:1930:CCF

[Loc30] S. B. Locke. A convenient container for fish specimens. Transactions of the American Fisheries Society, 60(1):250, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Locke:1932:DCS

[Loc32] S. B. Locke. Developments in the clean streams campaign. Transactions of the American Fisheries Society, 62(1):355–362, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Locke:1934:VCS

[Loc34] S. B. Locke. The value of clean streams. Transactions of the American Fisheries Society, 64(1):87–92, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lord:1930:RBS

[Lor30] Russell F. Lord. Rearing a brood stock of blackspotted trout. Transactions of the American Fisheries Society, 60(1):164–166, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lord:1932:NMG

[Lor32] Russell F. Lord. Notes on Montana graylings at the Pittsford, Vt., experimental trout hatchery. Transactions of the American Fisheries Society, 62(1):171–178, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lord:1933:TFT

[Lor33] Russell F. Lord. Type of food taken throughout the year by brook trout in a single Vermont stream with special reference to winter feeding. Transactions of the American Fisheries Society, 63(1):182–197, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lord:1934:HTF

[Lor34] Russell F. Lord. Hatchery trout as foragers and game fish. Transactions of the American Fisheries Society, 64(1):339–345, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lord:1935:THF

[Lor35] Russell F. Lord. The 1935 trout harvest from Furnace Brook, Vermont’s “test” stream. Transactions of the American Fisheries Society, 65(1):224–233, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leonard:1937:SMD

[LS37] J. W. Leonard and David S. Shetter. Studies on merganser depredations in Michigan trout waters. Transactions of the American Fisheries Society, 66(1):335–337, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

MacKay:1930:PPO

[Mac30a] H. H. MacKay. Pollution problems in Ontario. Transactions of the American Fisheries Society, 60(1):297–305, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

MacKay:1930:PSF

[Mac30b] H. H. MacKay. The present status of fish culture in the Province of Ontario. Transactions of the American Fisheries Society, 60(1):33–44, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Madsen:1937:PNF

[Mad37] David H. Madsen. Protection of native fishes in the national parks. Transactions of the American Fisheries Society, 66(1):395–397, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marston:1930:ETU

[Mar30] Alice T. Marston. The effect of temperature upon the ingestion of bacteria by the clam (Mya arenaria). Transactions of the American Fisheries Society, 60(1):269, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Markus:1932:EWT

[Mar32] Henry C. Markus. The extent to which temperature changes influence food consumption in large-mouth bass (Huro floridana). Transactions of the American Fisheries Society, 62(1):202–210, January 1932. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Markus:1933:ETU

[Mar33a] Henry C. Markus. The effects of tags upon fresh water fishes. Transactions of the American Fisheries Society, 63(1):319–325, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Markus:1933:PCF

[Mar33b] Henry C. Markus. The progress of the cooperative fish propagation at Rochester and Monroe County, New York. Transactions of the American Fisheries Society, 63(1):151–153, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Markus:1934:FOF

[Mar34] Henry C. Markus. The fate of our forage fish. Transactions of the American Fisheries Society, 64(1):93–96, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Markus:1935:ABT

[Mar35] Henry C. Markus. Americanized brown trout retain European characteristics. Transactions of the American Fisheries Society, 65(1): 258–260, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McKenzie:1930:RDF

[McK30] R. A. McKenzie. The reported decrease in fish life and the pollution of the Winnipeg River, Kenora, Ontario. Transactions of the American Fisheries Society, 60(1):311–323, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McLeod:1930:TSY

[McL30] A. W. McLeod. Thirty-six years’ experience in fish culture. Transactions of the American Fisheries Society, 60(1):45–47, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McNamara:1937:BFH

[McN37] Fred McNamara. Breeding and food habits of the pikes, (Esox lucius and esox vermiculatus). Transactions of the American Fisheries Society, 66(1):372–373, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meehean:1933:RFP

[Mee33] O. Lloyd Meehean. The role of fertilizers in pondfish. Transactions of the American Fisheries Society, 63(1):103–109, January 1933. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meehean:1934:RFP

[Mee34] O. Lloyd Meehean. The role of fertilizers in pondfish production II. Some ecological aspects. Transactions of the American Fisheries Society, 64(1):151–154, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meehean:1935:DFS

[Mee35] O. Lloyd Meehean. The dispersal of fertilizing substances in ponds. Transactions of the American Fisheries Society, 65(1):184–188, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mendez:1938:RFM

[Mén38] Santiago Rodríquez Méndez. The regulation of the fisheries in Mexico. Transactions of the American Fisheries Society, 67(1):41–45, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

MGonigle:1932:SET

[M’G32] R. H. M’Gonigle. Some effects of temperature, and their relation to fish culture. Transactions of the American Fisheries Society, 62(1): 119–125, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

MGonigle:1934:AFS

[M’G34] R. H. M’Gonigle. Algae, a factor in some hatchery mortalities. Transactions of the American Fisheries Society, 64(1):416–423, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:1934:PTS

[MGG^{+}34] Emmeline Moore, J. R. Greeley, C. W. Greene, H. M. Faigenbaum, F. R. Nevin, and H. K. Townes. A problem in trout stream management. Transactions of the American Fisheries Society, 64(1):68–80, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1932:OFB

[Mil32] Raymond Miller. Our fight for the black bass. Transactions of the American Fisheries Society, 62(1):363–365, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1937:MLS

[Mil37] Donald E. Miller. Michigan Lake Survey — a cooperative enterprise. Transactions of the American Fisheries Society, 66(1):128–130, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mobley:1931:CCC

[Mob31] Ben E. Mobley. The culture of channel cat fish (Ictaluras punctatus). Transactions of the American Fisheries Society, 61(1):171–173, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:1931:FPI

[Moo31] Emmeline Moore. A follow-up program on the impounded waters of the Sacandaga Reservoir. Transactions of the American Fisheries Society, 61(1):139–142, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:1932:CMS

[Moo32] Emmeline Moore. Certain minnows showing adaptability to conditions in impounded waters. Transactions of the American Fisheries Society, 62(1):290–291, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:1938:CCC

[Moo38] Emmeline Moore. A creel census at Chautauqua Lake, New York. Transactions of the American Fisheries Society, 67(1):130–138, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mottley:1932:PTK

[Mot32] C. McC. Mottley. The propagation of trout in the Kamloops District, British Columbia. Transactions of the American Fisheries Society, 62(1):144–151, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mottley:1933:SMR

[Mot33] C. McC. Mottley. The spawning migration of rainbow trout. Transactions of the American Fisheries Society, 63(1):80–84, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mottley:1934:ORR

[Mot34] C. McC. Mottley. The origin and relations of the rainbow trout. Transactions of the American Fisheries Society, 64(1):323–327, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mottley:1935:OPC

[Mot35] C. McC. Mottley. Objectives in the pond culture of salmonoid game fish. Transactions of the American Fisheries Society, 65(1):172–178, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mottley:1937:SEC

[Mot37] C. McC. Mottley. Sampling errors in calculating the weight of fish in ponds. Transactions of the American Fisheries Society, 66(1):422– 426, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mottley:1938:DFM

[Mot38a] C. McC. Mottley. Does the full moon affect rainbow trout fishing? Transactions of the American Fisheries Society, 67(1):212–214, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mottley:1938:LWR

[Mot38b] C. McC. Mottley. Loss of weight by rainbow trout at spawning time. Transactions of the American Fisheries Society, 67(1):207–210, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCay:1930:NRT

[MTC^{+}30] C. M. McCay, John W. Titcomb, Eben W. Cobb, Mary F. Crowell, and A. Tunison. The nutritional requirements of trout. Transactions of the American Fisheries Society, 60(1):127–139, January 1930. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCay:1931:NRT

[MTC^{+}31] C. M. McCay, A. Tunison, Mary Crowell, D. K. Tressler, S. P. Mac-Donald, John W. Titcomb, and Eben W. Cobb. The nutritional requirements of trout and chemical composition of the entire trout body. Transactions of the American Fisheries Society, 61(1):58–79, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

MacKay:1934:SOC

[MW34] H. H. MacKay and W. H. R. Werner. Some observations on the culture of maskinonge. Transactions of the American Fisheries Society, 64(1):

313–317, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Needham:1933:NUW

[NB33] P. R. Needham and L. A. Brown. Notes on the use of water fleas as fish food. Transactions of the American Fisheries Society, 63(1): 131–138, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Needham:1930:SSF

[Nee30] P. R. Needham. Studies on the seasonal food of brook trout. Transactions of the American Fisheries Society, 60(1):73–88, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Needham:1933:MSF

[Nee33] James G. Needham. Mayflies — a staple food of fishes in hill streams. Transactions of the American Fisheries Society, 63(1):178–181, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Needham:1934:QSS

[Nee34] P. R. Needham. Quantitative studies of stream bottom foods. Transactions of the American Fisheries Society, 64(1):238–247, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Needham:1935:CAF

[Nee35] P. R. Needham. A cheap, adjustable fish grader. Transactions of the American Fisheries Society, 65(1):313–315, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Needham:1938:NIS

[Nee38] Paul R. Needham. Notes on the introduction of Salmo nelsoni Evermann into California from Mexico. Transactions of the American Fisheries Society, 67(1):139–146, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nesbit:1933:NMM

[Nes33] Robert A. Nesbit. A new method of marking fish by means of internal tags. Transactions of the American Fisheries Society, 63(1):306–307, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Needham:1934:OSS

[NT34] P. R. Needham and A. C. Taft. Observations on the spawning of steelhead trout. Transactions of the American Fisheries Society, 64(1): 332–338, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nurnberger:1930:PAF

[Nur30] Patience Kidd Nurnberger. The plant and animal food of the fishes of Big Sandy Lake. Transactions of the American Fisheries Society, 60(1):253–259, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nurnberger:1931:OST

[Nur31] Patience Kidd Nurnberger. Observations on the spawning temperature of Luxilus cornutis. Transactions of the American Fisheries Society, 61(1):215, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Odell:1932:DDC

[Ode32] Theodore T. Odell. The depth distribution of certain species of fish in some of the lakes of New York. Transactions of the American Fisheries Society, 62(1):331–335, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Odell:1934:LHE

[Ode34] T. T. Odell. The life history and ecological relationships of the alewife (Pomolobus pseudoharengus — Wilson) in Seneca Lake, New York. Transactions of the American Fisheries Society, 64(1):118–126, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olson:1932:SOI

[Ols32] Theodore A. Olson. Some observations on the interrelationships of sunlight, aquatic plant life and fishes. Transactions of the American Fisheries Society, 62(1):278–289, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OMalley:1932:AFS

[O’M32] Henry O’Malley. The American Fisheries Society and the fisheries. Transactions of the American Fisheries Society, 62(1):351–354, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Osburn:1933:SIP

[Osb33] Raymond C. Osburn. Some important principles of fish conservation. Transactions of the American Fisheries Society, 63(1):91–97, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parnell:1934:FPT

[Par34] I. W. Parnell. Fish parasites and their importance. Transactions of the American Fisheries Society, 64(1):390–400, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pautzke:1938:SEC

[Pau38] Clarence F. Pautzke. Studies on the effect of coal washings on steelhead and cutthroat trout. Transactions of the American Fisheries Society, 67(1):232–233, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pentland:1930:CFD

[Pen30] Ernest S. Pentland. Controlling factors in the distribution of Gammarus. Transactions of the American Fisheries Society, 60(1):89–94, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phillips:1934:PDM

[Phi34] Mrs. J. T. Phillips. Plankton distribution in Manitou and Brome Lakes. Transactions of the American Fisheries Society, 64(1):461–463, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Powell:1931:MAR

[Pow31] A. M. Powell. Maryland’s activities in raising of trout and bass. Transactions of the American Fisheries Society, 61(1):80–82, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Powers:1938:FIS

[Pow38] Edwin B. Powers. Factors involved in the sudden mortality of fishes. Transactions of the American Fisheries Society, 67(1):271–281, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prevost:1934:CUP

[Pre34a] Gustave Prevost. A criticism of the use of potassium permanganate in fish culture. Transactions of the American Fisheries Society, 64(1): 304–306, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prevost:1934:NMA

[Pre34b] Gustave Prevost. A new method and apparatus for the accurate and rapid counting of fry and fingerlings. Transactions of the American Fisheries Society, 64(1):270–273, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prevost:1935:ESS

[Pre35] Gustave Prevost. Experimental stocking of speckled trout from the air. Transactions of the American Fisheries Society, 65(1):277–278, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pritchard:1932:RTP

[Pri32] A. L. Pritchard. Relation of tagging programs to the conservation of Pacific salmon off the Coast of British Columbia. Transactions of the American Fisheries Society, 62(1):88–93, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prytherch:1932:SMO

[Pry32] Herbert F. Prytherch. Scientific methods of oyster farming. Transactions of the American Fisheries Society, 62(1):72–75, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pujana:1938:UFR

[Puj38] Agustin Pujana. The undeveloped fisheries resources of Lower California. Transactions of the American Fisheries Society, 67(1):282–285, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Radebaugh:1934:EVT

[RA34] Gus H. Radebaugh and H. P. Kjerschow Agersborg. The economic value of treated sewage effluent in wildlife conservation, with special reference to fish and waterfowl. Transactions of the American Fisheries Society, 64(1):443–456, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Radcliffe:1931:PM

[Rad31] Lewis Radcliffe. Propagation of minnows. Transactions of the American Fisheries Society, 61(1):131–138, January 1931. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Radcliffe:1933:WFE

[Rad33] Lewis Radcliffe. Water-farming as exemplified by the oyster industry. Transactions of the American Fisheries Society, 63(1):154–157, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Radcliffe:1938:SFN

[Rad38] Lewis Radcliffe. Sponge farming: a necessity. Transactions of the American Fisheries Society, 67(1):46–51, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rawson:1932:PWL

[Raw32] D. S. Rawson. The pike of Waskesiu Lake, Saskatchewan: a preliminary report. Transactions of the American Fisheries Society, 62(1): 323–330, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rawson:1938:NRE

[Raw38] D. S. Rawson. Natural rearing enclosures for smallmouth black bass. Transactions of the American Fisheries Society, 67(1):96–104, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richardson:1934:MPO

[RB34] L. R. Richardson and G. W. Belknap. A method of planting Osmerus mordax Mitchill on a small scale. Transactions of the American Fisheries Society, 64(1):432–434, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Regan:1934:FBB

[Reg34] C. C. Regan. Floating bass-brooding equipment. Transactions of the American Fisheries Society, 64(1):143–145, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ricker:1930:FHS

[Ric30] William E. Ricker. Feeding habits of speckled trout in Ontario waters. Transactions of the American Fisheries Society, 60(1):64–72, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ricker:1932:UNF

[Ric32] W. E. Ricker. The utility of nets in fresh-water plankton investigations. Transactions of the American Fisheries Society, 62(1):292–303, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richardson:1934:OED

[Ric34] L. R. Richardson. Observations on the effects of dams on lakes and streams. Transactions of the American Fisheries Society, 64(1):457– 460, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richardson:1935:ROS

[Ric35] L. R. Richardson. A record of Octomitus salmonis Moore from Quebec. Transactions of the American Fisheries Society, 65(1):290–292, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richardson:1937:OPS

[Ric37] L. R. Richardson. Observations on the parasites of the speckled trout in Lake Edward, Quebec. Transactions of the American Fisheries Society, 66(1):343–356, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richardson:1938:OTI

[Ric38] Laurence R. Richardson. Observations on trichodinid infection (Cyclochaetosis) of Salvelinus fontinalis (Mitchill). Transactions of the American Fisheries Society, 67(1):228–231, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rivero:1937:ILB

[Riv37] Luis Howell Rivero. The introduced largemouth bass, a predator upon native Cuban fishes. Transactions of the American Fisheries Society, 66(1):367–368, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roach:1933:SPC

[Roa33] Lee S. Roach. Some physical, chemical, and biological studies of impounded waters in Ohio. Transactions of the American Fisheries Society, 63(1):265–270, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robertson:1937:HFG

[Rob37] Alexander Robertson. Hatching fry in gravel no. 2. Transactions of the American Fisheries Society, 66(1):279–282, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rodd:1930:UWA

[Rod30] J. A. Rodd. Unproductive water areas made productive. Transactions of the American Fisheries Society, 60(1):116–118, January 1930. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ross:1931:FFM

[Ros31] John H. Ross. The freshwater fishes of Missouri. Transactions of the American Fisheries Society, 61(1):119–120, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rushton:1937:FWF

[Rus37] W. Rushton. Fresh water fishery problems of the British Isles. Transactions of the American Fisheries Society, 66(1):383–391, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roach:1934:RAP

[RW34] Lee S. Roach and E. L. Wickliff. Relationship of aquatic plants to oxygen supply, and their bearing on fish life. Transactions of the American Fisheries Society, 64(1):370–378, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1932:OPF

[SA32] E. W. Surber and E. L. Aikens. Observations on plantings of fingerling rainbow, brown, and brook trout from circular pools. Transactions of the American Fisheries Society, 62(1):133–140, January 1932. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Salyer:1933:PSM

[Sal33] J. Clark Salyer. Predator studies in Michigan waters. Transactions of the American Fisheries Society, 63(1):229–239, January 1933. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Savage:1935:CIS

[Sav35a] James Savage. Copepod infection of speckled trout. Transactions of the American Fisheries Society, 65(1):334–339, January 1935. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Savage:1935:NC

[Sav35b] James Savage. Notes on costiasis. Transactions of the American Fisheries Society, 65(1):332–333, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Savage:1937:REG

[Sav37] James Savage. Report on the effect of gravel in raceways. Transactions of the American Fisheries Society, 66(1):283–286, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sc:1932:FCA

[Sc.32] B. W. Taylor M. Sc. Fish cultural activities of the Province of Quebec. Transactions of the American Fisheries Society, 62(1):224–227, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schultz:1935:SHC

[Sch35] Leonard P. Schultz. The spawning habits of the chub, Mylocheilus caurinus — a forage fish of some value. Transactions of the American Fisheries Society, 65(1):143–147, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schneberger:1937:BEI

[Sch37a] Edward Schneberger. The biological and economic importance of the smelt in Green Bay. Transactions of the American Fisheries Society, 66(1):139–142, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schultz:1937:SPB

[Sch37b] Leonard P. Schultz. Some principles of biological fisheries research. Transactions of the American Fisheries Society, 66(1):188–192, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scofield:1932:CFS

[Sco32] W. L. Scofield. Comments on fisheries statistics collected in California. Transactions of the American Fisheries Society, 62(1):94–99, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seale:1933:BSA

[Sea33] Alvin Seale. The brine shrimp (Artemia) as a satisfactory live food for fishes. Transactions of the American Fisheries Society, 63(1):129–130, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shetter:1937:MGR

[She37] David S. Shetter. Migration, growth rate, and population density of brook trout in the North Branch of the Au Sable River, Michigan. Transactions of the American Fisheries Society, 66(1):203–210, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shirts:1931:PBB

[Shi31] Walter Shirts. The propagation of black bass in Indiana. Transactions of the American Fisheries Society, 61(1):116–118, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shillington:1934:SBS

[Shi34] K. G. Shillington. Selective breeding of speckled trout. Transactions of the American Fisheries Society, 64(1):274–275, January 1934. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stevens:1931:FI

[SK31] Margaret L. Stevens and W. M. Keil. Furunculosis investigations. Transactions of the American Fisheries Society, 61(1):193–198, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1931:LCA

[SM31] E. W. Surber and O. L. Meehean. Lethal concentrations of arsenic for certain aquatic organisms. Transactions of the American Fisheries Society, 61(1):225–239, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1932:FWE

[Smi32] M. W. Smith. The fertilization of water in experimental ponds. Transactions of the American Fisheries Society, 62(1):317–322, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1933:CSD

[Smi33] Osgood R. Smith. The Caledonia shrimp dikerogammarus. Transactions of the American Fisheries Society, 63(1):120–128, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1934:DOC

[Smi34] M. W. Smith. The dissolved oxygen content of fertilized waters. Transactions of the American Fisheries Society, 64(1):408–415, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1935:PNF

[Smi35a] M. W. Smith. A preliminary note on the fish population of Lake Jesse, Nova Scotia. Transactions of the American Fisheries Society, 65(1): 297–299, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1935:UCS

[Smi35b] M. W. Smith. The use of copper sulphate for eradicating the predatory fish population of a lake. Transactions of the American Fisheries Society, 65(1):101–114, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1935:BHS

[Smi35c] Osgood R. Smith. The breeding habits of the stone roller minnow (Campostoma anomalum Rafinesque). Transactions of the American Fisheries Society, 65(1):148–151, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1937:ONV

[Smi37a] Lester Smith. Observations on natural versus artificial propagation of commercial species of fish in the Great Lakes region. Transactions of the American Fisheries Society, 66(1):56–62, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1937:NFD

[Smi37b] M. W. Smith. Notes on the food of Daphnia pulex (De Geer) in fertilized water. Transactions of the American Fisheries Society, 66(1): 287–290, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1937:CCD

[Smi37c] Osgood R. Smith. A creel census from Dry River, New Hampshire. Transactions of the American Fisheries Society, 66(1):313–315, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1938:PAF

[Smi38] M. W. Smith. A preliminary account of the fish populations in certain Nova Scotian lakes. Transactions of the American Fisheries Society, 67(1):178–183, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snyder:1932:TSB

[Sny32] J. P. Snyder. Tagged smallmouth black bass in Lake Ontario, n. y. Transactions of the American Fisheries Society, 62(1):380–381, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1937:SOM

[SO37] Thaddeus Surber and Theodore A. Olson. Some observations on Minnesota fish ponds. Transactions of the American Fisheries Society, 66(1):104–127, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spier:1934:DSA

[Spi34] Miss J. D. Spier. The dominating species of aquatic plants in Brome Lake and Lake Manitou (Province of Quebec). Transactions of the American Fisheries Society, 64(1):464–468, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stone:1938:GHF

[Sto38] Udell Bennett Stone. Growth, habits, and fecundity of the ciscoes of Irondequoit Bay, New York. Transactions of the American Fisheries Society, 67(1):234–245, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1930:QMS

[Sur30a] E. W. Surber. A quantitative method of studying the food of small fishes. Transactions of the American Fisheries Society, 60(1):158–163, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1930:MQB

[Sur30b] Eugene W. Surber. A method of quantitative bottom fauna and facultative plankton study employed in a year’s study of slough biology. Transactions of the American Fisheries Society, 60(1):187–198, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1931:SAC

[Sur31a] Eugene W. Surber. Sodium arsenite for controlling submerged vegetation in fish ponds. Transactions of the American Fisheries Society, 61(1):143–149, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1931:FCS

[Sur31b] Thaddeus Surber. Fish cultural successes and failures in Minnesota. Transactions of the American Fisheries Society, 61(1):240–246, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1933:OCP

[Sur33a] Eugene W. Surber. Observations on circular pool management. Transactions of the American Fisheries Society, 63(1):139–143, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1933:QSR

[Sur33b] Eugene W. Surber. A quantitative study of rainbow trout production in one mile of stream. Transactions of the American Fisheries Society, 63(1):251–256, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1933:RLT

[Sur33c] Thaddeus Surber. Rearing lake trout to maturity. Transactions of the American Fisheries Society, 63(1):64–68, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1935:ECD

[Sur35a] Eugene W. Surber. Effects of carbon dioxide on the development of trout eggs. Transactions of the American Fisheries Society, 65(1): 194–203, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1935:TFE

[Sur35b] Eugene W. Surber. Trout feeding experiments with natural food (Gammarus fasciatus). Transactions of the American Fisheries Society, 65(1):300–304, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1937:RTB

[Sur37] Eugene W. Surber. Rainbow trout and bottom fauna production in one mile of stream. Transactions of the American Fisheries Society, 66(1):193–202, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Taft:1934:CSE

[Taf34] A. C. Taft. California steelhead experiments. Transactions of the American Fisheries Society, 64(1):248–251, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tarzwell:1931:TSI

[Tar31] Clarence M. Tarzwell. Trout stream improvement in Michigan. Transactions of the American Fisheries Society, 61(1):48–57, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tarzwell:1937:EEV

[Tar37] Clarence M. Tarzwell. Experimental evidence on the value of trout stream improvement in Michigan. Transactions of the American Fisheries Society, 66(1):177–187, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tarzwell:1938:FIF

[Tar38] Clarence M. Tarzwell. Factors influencing fish food and fish production in southwestern streams. Transactions of the American Fisheries Society, 67(1):246–255, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Taylor:1935:CSW

[Tay35] B. W. Taylor. A cheap supply of warm water for winter hatching and feeding of speckled trout. Transactions of the American Fisheries Society, 65(1):340–341, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tester:1930:SHS

[Tes30] Albert L. Tester. Spawning habits of the small-mouthed black bass in Ontario waters. Transactions of the American Fisheries Society, 60 (1):53–61, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tester:1933:PHC

[Tes33] Albert L. Tester. Populations of herring in the coastal waters of British Columbia. Transactions of the American Fisheries Society, 63(1): 286–291, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trautman:1935:WDP

[TH35] Milton B. Trautman and Carl L. Hubbs. When do pike shed their teeth? Transactions of the American Fisheries Society, 65(1):261– 266, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tidd:1934:RIG

[Tid34] Wilbur M. Tidd. Recent infestations of goldfish and carp by the “anchor parasite”, Lernaea carassii. Transactions of the American Fisheries Society, 64(1):176–180, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Titcomb:1930:MAF

[Tit30] John W. Titcomb. More about the fresh-water shrimp. Transactions of the American Fisheries Society, 60(1):95–100, January 1930. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Taylor:1934:TLG

[TL34] B. W. Taylor and R. C. Lindsay. Trout lakes in Gaspé County. Transactions of the American Fisheries Society, 64(1):424–431, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tunison:1933:NRB

[TM33] A. V. Tunison and C. M. McCay. The nutritional requirements of brook trout. Transactions of the American Fisheries Society, 63(1): 167–177, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tunison:1935:NRT

[TM35] A. V. Tunison and C. M. McCay. The nutritional requirements of trout. Transactions of the American Fisheries Society, 65(1):359–375, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trautman:1932:PVS

[Tra32] Milton B. Trautman. The practical value of scientific data in a stocking policy for Ohio. Transactions of the American Fisheries Society, 62(1):258–260, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trautman:1933:GEP

[Tra33] Milton B. Trautman. The general effects of pollution on Ohio fish life. Transactions of the American Fisheries Society, 63(1):69–72, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Truitt:1932:SFW

[Tru32] R. V. Truitt. Scientific fisheries work in Maryland. Transactions of the American Fisheries Society, 62(1):50–56, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Truitt:1937:ISF

[TV37a] R. V. Truitt and V. D. Vladykov. The importance of sport fishing in Maryland. Transactions of the American Fisheries Society, 66(1): 403–405, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Truitt:1937:SBI

[TV37b] R. V. Truitt and V. D. Vladykov. Striped bass investigations in the Chesapeake Bay. Transactions of the American Fisheries Society, 66(1):225–226, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanOosten:1930:DLE

[Van30] John Van Oosten. The disappearance of the Lake Erie cisco — a preliminary report. Transactions of the American Fisheries Society, 60(1):204–214, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanOosten:1932:EMT

[Van32] John Van Oosten. Experiments on the mesh of trapnets and legislation of the commercial fisheries of Lake Erie. Transactions of the American Fisheries Society, 62(1):100–107, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanOosten:1934:VQC

[Van34] John Van Oosten. The value of questionnaires in commercial fisheries regulations and surveys. Transactions of the American Fisheries Society, 64(1):107–117, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanOosten:1935:LJD

[Van35] John Van Oosten. Logically justified deductions concerning the Great Lakes fisheries exploded by scientific research. Transactions of the American Fisheries Society, 65(1):71–75, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanDerSchalie:1937:RGP

[Van37a] Henry Van Der Schalie. The relationship of the gravid periods of certain mussels in Michigan to the pearl button industry. Transactions of the American Fisheries Society, 66(1):406–410, January 1937. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanOosten:1937:DSO

[Van37b] John Van Oosten. The dispersal of smelt, Osmerus mordax (Mitchill), in the Great Lakes region. Transactions of the American Fisheries Society, 66(1):160–171, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanOosten:1937:GLF

[Van37c] John Van Oosten. The Great Lakes fisheries: Their proper management for sustained yields. Transactions of the American Fisheries Society, 66(1):131–138, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanOosten:1938:FLT

[VD38] John Van Oosten and Hilary J. Deason. The food of the lake trout (Cristivomer namaycush Namaycush) and of the lawyer (Lota maculosa) of Lake Michigan. Transactions of the American Fisheries Society, 67(1):155–177, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Viosca:1931:PBR

[Vio31a] Percy Viosca, Jr. Principles of bullfrog (Rana catesbiana) culture. Transactions of the American Fisheries Society, 61(1):262–269, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Viosca:1931:SSM

[Vio31b] Percy Viosca, Jr. The southern small-mouth black bass: Micropterus pseudaplites Hubbs. Transactions of the American Fisheries Society, 61(1):95–105, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Viosca:1935:SPI

[Vio35] Percy Viosca, Jr. Statistics on the productivity of inland waters: The master key to better fish culture. Transactions of the American Fisheries Society, 65(1):350–358, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

vonIhering:1935:FIN

[vIW35] Rodolpho von Ihering and Stillman Wright. Fisheries investigations in Northeast Brazil. Transactions of the American Fisheries Society, 65(1):267–271, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vladykov:1938:SBR

[VW38a] V. D. Vladykov and D. H. Wallace. Is the striped bass (Roccus lineatus) of Chesapeake Bay a migratory fish? Transactions of the American Fisheries Society, 67(1):67–86, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vladykov:1938:RPS

[VW38b] V. D. Vladykov and D. H. Wallace. Remarks on populations of the shad (Alosa sapidissima) along the Atlantic Coast region. Transactions of the American Fisheries Society, 67(1):52–66, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wales:1935:FTC

[Wal35] Joseph H. Wales. The feeding of trout in California hatcheries. Transactions of the American Fisheries Society, 65(1):305–308, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ward:1931:NFP

[War31] Fern Ward. Notes on the food and parasites of the mosquito fish (Gambusia holbrooki) in Florida. Transactions of the American Fisheries Society, 61(1):208–214, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Watt:1932:LFF

[Wat32] A. M. Watt. The Law of Fish and Fishing in the Province of Quebec. Transactions of the American Fisheries Society, 62(1):384–385, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wiebe:1934:PSP

[WBF34] A. H. Wiebe, J. G. Burr, and H. E. Faubion. The problem of stream pollution in Texas with special reference to salt water from the oil fields. Transactions of the American Fisheries Society, 64(1):81–86, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wynne-Edwards:1932:BHB

[WE32] V. C. Wynne-Edwards. The breeding habits of the black-headed minnow (Pimephales promelas Raf.). Transactions of the American Fisheries Society, 62(1):382–383, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Webster:1931:SF

[Web31] B. O. Webster. A successful fishway. Transactions of the American Fisheries Society, 61(1):247–257, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Werner:1934:NUP

[Wer34] W. H. R. Werner. A note on the use of physiological saline as defined herein as an aid in the artificial spawning of speckled trout. Transactions of the American Fisheries Society, 64(1):346–350, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Westerman:1932:DWT

[Wes32] Fred A. Westerman. The deep water trap net and its relation to the Great Lakes fisheries. Transactions of the American Fisheries Society, 62(1):64–71, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Westerman:1934:TAM

[Wes34] Fred A. Westerman. Transactions of the 64th Annual Meeting of the American Fisheries Society: Montreal, Canada, September 12, 13 and 14, 1934. Transactions of the American Fisheries Society, 64(1):13–46, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Westerman:1938:ENF

[Wes38] Fred A. Westerman. Exploring new fields in fisheries management. Transactions of the American Fisheries Society, 67(1):105–113, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wiebe:1933:OCL

[WF33] A. H. Wiebe and Arno C. Fuller. The oxygen consumption of largemouth black bass (Huro floridana) fingerling. Transactions of the American Fisheries Society, 63(1):208–214, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Watson:1935:NVB

[WF35] Vernon K. Watson and Carl R. Fellers. Nutritive value of the blue crab (Callinectes sapidus), and sand crab (Platyonichus ocellatus Latreille). Transactions of the American Fisheries Society, 65(1):342– 349, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wiebe:1931:ACL

[WGS31] A. H. Wiebe, E. G. Gross, and D. H. Slaughter. The arsenic content of large-mouth black bass (Micropterus salmoides Lacépède) fingerlings. Transactions of the American Fisheries Society, 61(1):150–163, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

White:1930:SOE

[Whi30] H. C. White. Some observations on the Eastern brook trout (S. Fontinalis) of Prince Edward Island. Transactions of the American Fisheries Society, 60(1):101–108, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

White:1934:SFT

[Whi34a] H. C. White. Some facts and theories concerning the Atlantic salmon. Transactions of the American Fisheries Society, 64(1):360–362, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

White:1934:SPB

[Whi34b] H. C. White. The spawning period of brook trout, S. Fontinalis. Transactions of the American Fisheries Society, 64(1):356–359, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wicks:1930:PUM

[Wic30] Judson L. Wicks. Pollution of the Upper Mississippi River. Transactions of the American Fisheries Society, 60(1):286–296, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wickliff:1931:FRO

[Wic31] E. L. Wickliff. Fishery research by the Ohio Division of Conservation. Transactions of the American Fisheries Society, 61(1):199–207, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wickliff:1932:NIW

[Wic32] E. L. Wickliff. Are newly impounded waters in Ohio suitable for fish life? Transactions of the American Fisheries Society, 62(1):275–277, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wickliff:1933:PVD

[Wic33a] E. L. Wickliff. The practical value of determining the fertility of whitefish eggs. Transactions of the American Fisheries Society, 63(1): 144–150, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wickliff:1933:RFT

[Wic33b] E. L. Wickliff. Returns from fish tagged in Ohio. Transactions of the American Fisheries Society, 63(1):326–331, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wickliff:1933:SFR

[Wic33c] E. L. Wickliff. A summary of fisheries research. Transactions of the American Fisheries Society, 63(1):257–264, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wickliff:1935:TAM

[Wic35] E. L. Wickliff. Transactions of the 65th Annual Meeting of the American Fisheries Society: Tulsa, Oklahoma, September 9, 10 and 11, 1935. Transactions of the American Fisheries Society, 65(1):13–47, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wickliff:1938:ARF

[Wic38] Edward L. Wickliff. Additional returns from fish tagged in Ohio. Transactions of the American Fisheries Society, 67(1):211, January 1938. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wiebe:1930:NEY

[Wie30] A. H. Wiebe. Notes on the exposure of young fish to varying concentrations of arsenic. Transactions of the American Fisheries Society, 60(1):270–278, January 1930. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wiebe:1931:NES

[Wie31] A. H. Wiebe. Notes in the exposure of several species of fish to sudden changes in the hydrogenion concentration of the water and to an atmosphere of pure oxygen. Transactions of the American Fisheries Society, 61(1):216–224, January 1931. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wiebe:1934:NDD

[Wie34] A. H. Wiebe. Nocturnal depressions in the dissolved oxygen in fish ponds with special reference to an excess of coarse vegetation and of fertilizers (Texas). Transactions of the American Fisheries Society, 64(1):181–188, January 1934. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wiebe:1935:ECO

[Wie35] A. H. Wiebe. The effect of crude oil on fresh-water fish. Transactions of the American Fisheries Society, 65(1):324–331, January 1935. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weymouth:1932:SLH

[WLA32] Frank W. Weymouth, Milton J. Lindner, and W. W. Anderson. A summary of the life history of the common shrimp (Penaeus setiferus) of the South Atlantic and Gulf Coasts of the United States. Transactions of the American Fisheries Society, 62(1):108–110, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wiebe:1932:ASS

[WM32] A. H. Wiebe and Alfred M. McGavock. The ability of several species of fish to survive on prolonged exposure to abnormally high concentrations of dissolved oxygen. Transactions of the American Fisheries Society, 62(1):267–274, January 1932. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wolf:1935:UPP

[Wol35] Louis E. Wolf. The use of potassium permanganate in the control of fish parasites. Transactions of the American Fisheries Society, 65(1): 88–100, January 1935. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wolf:1937:ABD

[Wol37a] Louis E. Wolf. An air bladder disease in lake trout fingerlings. Transactions of the American Fisheries Society, 66(1):359–363, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wolf:1937:BCG

[Wol37b] Louis E. Wolf. Blood clots in the gills of trout. Transactions of the American Fisheries Society, 66(1):369–371, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wickliff:1937:SSI

[WR37] E. L. Wickliff and Lee S. Roach. Some studies of impounded waters in Ohio. Transactions of the American Fisheries Society, 66(1):76–86, January 1937. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wright:1933:SLI

[WT33] Stillman Wright and Wilbur M. Tidd. Summary of limnological investigations in Western Lake Erie in 1929 and 1930. Transactions of the American Fisheries Society, 63(1):271–285, January 1933. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).